### v1

# Enhancing Robustness of CLIP to Common Corruptions through Bimodal Test-Time Adaptation

Abstract
Although open-vocabulary classification models like Contrastive Language Image Pretraining (CLIP) have demonstrated strong zero-shot learning capabilities, their robustness to common image corruptions remains poorly understood. Through extensive experiments, we show that zero-shot CLIP lacks robustness to common image corruptions at increasing severity levels during test-time, necessitating the adaptation of CLIP to unlabeled corrupted images using test-time adaptation (TTA). However, we found that existing TTA methods have severe limitations in adapting CLIP due to their unimodal nature. To address these limitations, we propose BAT-CLIP, a bimodal TTA method specially designed to improve CLIP’s robustness to common image corruptions. The key insight of our approach is not only to adapt the visual encoders for better image feature extraction but also to strengthen the alignment between image and text features by promoting a stronger association between the image class prototype, computed using pseudo-labels, and the corresponding text feature. We evaluate our approach on benchmark image corruption datasets and achieve state-of-the-art results in TTA for CLIP, specifically for domains involving image corruption. Particularly, with a ViT-B/16 vision backbone, we obtain mean accuracy improvements of 9.7%, 5.94%, and 5.12% for CIFAR-10C, CIFAR-100C, and ImageNet-C, respectively.
1 Introduction
(a) TPT [38]
(b) VTE [9]
(c) BAT-CLIP (Ours)
(d) Performance comparison
Figure 1: Comparison of BAT-CLIP with other online TTA approaches using CLIP: TPT [38], and VTE [9]. a) TPT optimizes text prompts only for a single test image, making it unimodal. b) VTE considers an ensemble of prompts without model updates. Both methods consider the generation of multiple augmentations of the test image. c) BAT-CLIP is a bimodal approach, that adapts the LayerNorm parameters of the vision and text encoders, maximizing alignment between class prototypes and text features while increasing the inter-class separability of prototypes.
The emergence of large pre-trained vision-language models (VLMs), such as CLIP [31], has led to their widespread adoption in various visual recognition tasks, including segmentation [20, 25], detection [2, 23], classification [52, 51], and image generation [45, 32, 33]. Thanks to supervision from massive corpora of paired language and image data, VLMs like CLIP demonstrate strong zero-shot capabilities for these downstream tasks.
Despite CLIP’s successes in such important applications, its robustness when faced with corrupted images remains largely underexplored. Our motivation stems from the fact that the vision perception system of humans exhibits a level of robustness that real-world vision systems are yet to achieve. For example, models deployed for safety-critical applications like autonomous driving [1], could face rapid distributional shifts of blurriness, pixel changes, snowy nights, or other weather conditions [35]. For instance, our findings on the zero-shot performance of CLIP with a ResNet-101 [14] vision backbone reveals that the accuracy on the test set of CIFAR100 [17] with Gaussian noise of severity level 5, plummets to 10.79% from 49% on the clean set. Similar trends are observed with ViT-B/16, -B/32, and -L/14 [10] as backbones. Such behaviors can result in significant performance degradation of models when encountering image corruption in real-world scenarios, particularly in applications like self-driving cars.
This challenge is not unique to CLIP; it reflects a broader issue in computer vision tasks [7, 11, 12, 5], which often rely on the assumption that training and test data share the same distribution. When test distributions differ, model adaptation becomes essential to accommodate this shift and maintain accurate predictions. In response to this issue, test-time adaptation (TTA) has garnered significant interest [46, 8]. TTA aims to adapt a pre-trained model to unlabeled batches of test data from varying domains in an online fashion—updating the model for each test batch without access to the source dataset. Several TTA methods proposed in the literature have been effective in mitigating domain shifts [46, 36, 34, 28, 42, 3]. In this paper, we use the terms distributions and domains interchangeably.
While there have been few new works on using CLIP for online TTA [38, 9], they come with certain limitations. For example, test-time prompt tuning (TPT) [38] tunes the text prompts on the text encoder alone and generates multiple random augmented views, for each test image. The text prompts, initialized to pre-trained values, are optimized by minimizing the marginal entropy of the confident model predictions. The prompts are reset after adaptation to each image. However, such a method is expensive and slow due to performing multiple forward passes through the vision encoder of CLIP, for each image. Also, it relies on hand-crafted prompts for initialization, making it impractical at test-time. Another recent approach, Vision-Text-Space Ensemble (VTE) [9], uses an ensemble of different prompts as input to CLIP’s text encoder while keeping both the encoders frozen. Since the vision encoder is frozen, it struggles to adapt images with severe noise effectively.
While TPT effectively improves CLIP’s test generalization by dynamically tuning the text prompts, it remains primarily a unimodal approach. This limits the capacity of a multimodal model like CLIP to fully leverage its multimodal nature for adaptation. Specifically, it prevents the encoders from jointly adjusting their features, resulting in suboptimal alignment between the visual and text modalities. During TTA, there is no transfer of knowledge, via gradients, between the visual and text encoders, as the adaptation process focuses solely on one modality. For instance, when a test image includes common corruptions, the text prompts adapt, but the vision encoder features remain fixed. As a result, the learned prompts lack awareness of the test image distribution, leading to a less effective adaptation.
To address these core limitations, we propose, BAT-CLIP, a bimodal adaptation approach of CLIP for online TTA, where both the visual and text encoders are jointly adapted by exploiting CLIP’s shared feature space of images and text. The overall objective is to achieve a strong alignment between image features and text features to enable more effective multimodal learning and adaptation. The adaptation procedure is two-fold: 1) Inspired by [47] for efficient fine-tuning, we adapt the norm layers i.e., LayerNorm parameters of both encoders. However, such a model update does not consider the alignment of the encoder features. So, to improve the alignment between class-specific visual and text features, we introduce a projection matching loss that maximizes the projection of the visual class prototypes with their corresponding text features. 2) To learn more discriminative visual features, we increase the cosine distance between the class prototypes, promoting a more distinct separation in the image feature space. Furthermore, our method is general-purpose, as it does not rely on hand-crafted prompt templates or their ensembles, unlike VTE. We leverage batches of test samples for TTA, rather than focusing on single-image adaptation, which makes our method fast. In the Supplementary, we show several examples of classification results, as a comparison.
Our novelty lies in performing a bimodal online test-time adaptation (TTA) of CLIP for classification tasks, to improve CLIP’s robustness. We draw a comparison of our method against TPT and VTE in Figure 1. Our main contributions are as follows:
  • We start with an in-depth analysis of CLIP’s zero-shot performance across different visual backbones on common image corruptions, at test-time, for varying severity levels. While CLIP shows strong performance on clean images, its performance drops notably on corrupted images.
  • To address the unimodal limitations highlighted earlier, our proposed bimodal adaptation of CLIP encoders at test-time aims to enhance alignment by maximizing the projection of class-wise prototypes onto their corresponding text features. Simultaneously, we increase the cosine distance between class prototypes to encourage the learning of more discriminative features, making the test adaptation process more flexible and robust.
  • We conduct extensive experiments benchmarking our method against prior TTA approaches using CLIP, i.e., TPT and VTE, for online test-time adaptation. We also compare BAT-CLIP against four TTA baselines, adopted for CLIP. Our approach achieves state-of-the-art performance for CLIP adaptation to common corruptions.
2 Related Works
Online Test-Time Adaptation (TTA). The objective of TTA is to adapt a pre-trained model, with no access to the source data, to incoming batches of unlabelled test data of a specific domain [46, 42, 36, 34, 28, 3, 49, 6]. In this approach, the model is reset to its pre-trained state after adapting to each target domain. As the updates are performed online, TENT [46] updates the affine parameters of the normalization layers and minimizes the entropy [37] of the model predictions. RoTTA [47] proposes a robust feature normalization method, leveraging a memory bank to track test data distribution and a teacher-student framework for adaptive reweighting and updates. RPL [34] argues that self-learning during adaptation via entropy minimization and pseudo-labels is beneficial. They propose the usage of a generalized cross-entropy loss for adaptation. SAR [28] filters noisy test samples that cause a performance drop identified from the gradient space. However, none of these works utilize CLIP for TTA.
TTA using CLIP. Lately, CLIP has been finding extensive applications for TTA. TPT [38] was the first work to propose prompt tuning using CLIP at test-time in an online manner. However, this method is computationally intensive for generating multiple views, per image. Additionally, a key limitation of this approach is its reliance on using CLIP’s default hand-crafted template for prompt initialization, with resets after adaptation to each image. A similar line of work was done in VTE [9] where ensembles are created in the text and vision space without any CLIP parameter update. [40] focus on single-image TTA for out-of-distribution detection and test generalization [22, 19]. WATT [29] borrows the idea of “test-time training” i.e., over several optimization steps, it adapts CLIP’s vision encoder using multiple text prompt templates, for a batch of test images. However, this TTA setting differs from ours where we update the model only once on a single test batch. SwapPrompt [26] explores the idea of self-supervised contrastive learning for prompt learning. HisTPT [48] relies on a memory bank, to store features of recent test samples, for prompt learning. Such a setup is impractical in real-time and is expensive. To summarize — while TPT is unimodal, our approach is bimodal, involving on-the-fly joint optimization of both encoders for test adaptation, leading to a strong multimodal alignment between features. We use a generic prompt template, avoiding prompt engineering and enabling real-time deployment.
3 Zero-shot performance analysis of CLIP to common image corruptions
While CLIP generalizes well to new concepts across vision-language modalities [4, 13], its performance under image corruptions is less explored. This section evaluates CLIP’s zero-shot capabilities in real-world scenarios with domain shifts caused by common corruptions, focusing on two primary areas: zero-shot performance and the need for adaptation to address domain shifts effectively.
Vision Backbones. We evaluate the robustness with a ResNet-101 (RN101) vision backbone [14] and three Vision Transformer backbones (ViT-B/16, ViT-B/32, ViT-L/14) [10]. The models are tested on their zero-shot classification performance i.e., accuracy.
Datasets. For all the experiments in this paper, we utilize the CIFAR-10C, CIFAR-100C, and ImageNet-C [15] datasets, which are standard benchmark datasets to evaluate the robustness of classification models. Each dataset contains 15 distinct corruption types as tasks (e.g., Gaussian noise, Shot noise, …). Each corruption is applied at 5 different severity levels to the test sets of CIFAR10, CIFAR100 [17], and ImageNet [7], acting as source test sets, and allowing us to systematically evaluate the model’s performance under increasing degrees of image degradation.
[…]
$p(y=c|x_{k,i}^{t})$ $=\frac{\text{exp}(\text{sim}(v_{k,i}^{t},z_{c})/\tau)}{\sum_{j}{}\text{exp}(\text{sim}(v_{k,i}^{t},z_{j})/\tau)}$ (1)
$\text{sim}(v,z)$ $=\frac{v^{T}\cdot z}{||v||_{2}\cdot||z||_{2}}$
where $\text{sim}(\cdot)$ is the cosine similarity and $\tau$ is the softmax temperature from CLIP’s pre-training stage. The text features $\mathcal{Z}$, in zero-shot evaluation, are always pre-computed. We also draw a comparison with CLIP performance on respective source test sets and follow  [^1] https://github.com/LAION-AI/CLIP_benchmark for this implementation. Throughout this paper, the same TTA problem setup is employed.
Figure 2: Task-wise mean accuracy (%) of zero-shot CLIP across corruption severity levels. [Top]: ResNet-101 backbone. [Bottom]: ViT-B/16 backbone. Dashed lines show zero-shot CLIP performance on corresponding source test sets.
Prompt Template Backbone
CIFAR-10C
CIFAR-100C
ImageNet-C
“a photo of a ¡CLS¿.”
RN101
42.30 (-36.50) 20.62 (-25.48) 14.97 (-46.23)
ViT-B/16
61.16 (-28.84) 35.78 (-30.82) 24.51 (-42.16)
ViT-B/32
59.00 (-29.20) 31.79 (-30.51) 23.20 (-38.80)
ViT-L/14
75.84 (-19.36) 47.82 (-27.78) 39.55 (-33.95)
“a bad photo of a ¡CLS¿.”
RN101
42.73 (-36.47) 20.18 (-27.02) 12.87 (-49.43)
ViT-B/16
62.97 (-28.03) 36.37 (-30.33) 25.84 (-42.26)
ViT-B/32
60.25 (-29.75) 31.30 (-31.00) 23.58 (-39.22)
ViT-L/14
76.08 (-19.22) 48.09 (-25.41) 39.37 (-34.73)
“a blurry photo of a ¡CLS¿.”
RN101
45.32 (-34.98) 20.28 (-26.12) 15.80 (-45.70)
ViT-B/16
62.49 (-28.31) 35.17 (-30.83) 25.27 (-42.03)
ViT-B/32
57.85 (-30.65) 31.47 (-30.85) 23.40 (-38.80)
ViT-L/14
73.77 (-20.73) 48.07 (-25.73) 39.22 (-34.18)
“a noisy photo of a ¡CLS¿.”
RN101
44.56 (-34.34) 20.71 (-25.89) 15.48 (-45.72)
ViT-B/16
63.03 (-27.87) 35.14 (-30.76) 25.05 (-41.35)
ViT-B/32
59.76 (-29.24) 31.53 (-30.37) 23.11 (-38.19)
ViT-L/14
75.36 (-19.54) 47.87 (-27.63) 38.78 (-34.42)
Table 1: Mean classification accuracy (%) across all corruption tasks with various prompt templates. We report the accuracy drop (-) (%) relative to performance on the corresponding source test set.
3.1 Sensitivity of CLIP to image corruption severity
We analyze CLIP’s zero-shot performance by progressively increasing corruption severity for RN101 and ViT-B/16 vision backbones. We report mean accuracy and overall performance in Figure 2. Despite CLIP’s robust multimodal feature space, accuracy drops significantly with an increase in corruption severity, regardless of the backbone. Specifically, for CIFAR-10C with a ViT-B/16 backbone, we observe accuracy as low as 37.92%, at a severity level of 5, for Gaussian noise. Similarly, for CIFAR-100C and ImageNet-C, the mean accuracy rates are as low as 35.79% and 24.51% at a severity level of 5, respectively. Results on additional vision backbones are in the Supplementary.
Analysis. CLIP’s zero-shot classification accuracy varies across models and datasets. Comparing the results to source test sets, for CIFAR10, RN101 achieves 78.8%, improving to 90.1% with ViT-B/16. On CIFAR100, accuracies are 46.1% and 66.6%, respectively. For ImageNet, RN101 scores 61.2%, while ViT-B/16 reaches 67.7%. More importantly, the key takeaway from Figure 2 is — even a slight increase in severity to level 1, for a majority of the corruption tasks, leads to a noticeable drop in accuracy. One plausible explanation for CLIP’s subpar performance is that the parameters of $f_{vis}$ were not optimized for such corruptions during pre-training. In zero-shot classification, the visual features from a given domain may lack the robustness and richness necessary to align well with their corresponding text features. This results in lower likelihoods and thus, higher misclassification rates.
3.2 Sensitivity of CLIP to prompt templates
In this analysis, we evaluate the impact of prompt engineering by providing “relevant” prompt templates to $f_{txt}$ for TTA. Each prompt adds context to help CLIP extract more relevant text features. We report the mean classification accuracy (in %) across RN101, ViT-B/16, ViT-B/32, and ViT-L/14 backbones at an image corruption severity level of 5 for all datasets, with results summarized in Table 1. The absolute accuracy drop compared to zero-shot performance on the source test set is also reported.
Analysis. It is interesting to observe that, irrespective of the backbones used, we do not see any drastic changes in the mean accuracy, for different “relevant” prompt templates. However, the major concern arises in the performance gap of each model and the zero-shot CLIP performance on the corresponding source test set, for the same prompt template. This discrepancy highlights the limited robustness of CLIP’s text encoder $f_{txt}$ to prompt selection in the context of image corruption. A key reason for this is that, despite the use of “relevant” prompts, the text and visual features remain largely independent and unaware of one another.
As expected, RN101 performs worse than the ViT-based backbones, primarily due to its lack of global attention-based modeling inherent to transformers [44]. Therefore, for the remainder of the experiments in this paper, we focus on ViT-based backbones, specifically ViT-B/16 and ViT-B/32. Due to GPU memory limitations, we do not use ViT-L/14 for model adaptation.
Unsuitability of prompt template selection at test-time. Additionally, at test-time, it is impractical to perform prompt engineering or optimize prompt vectors since 1) Choosing different prompt templates for generating text features is extremely tedious and time-consuming. 2) In real-time deployment involving prompt-tuning, such prompts cannot quickly estimate the distribution of incoming test batches. TPT [38] optimizes pre-trained text prompts for each test image, turning out to be suboptimal since the prompts are optimized ignorant of the distribution of the test image.
4 BAT-CLIP
Figure 3: BAT-CLIP not only adapts the visual encoder for highly discriminative image features but also promotes a strong alignment between image and text features by adapting the text encoder too, leading to improved performance following test-time adaptation. We adapt only the LayerNorm parameters of CLIP encoders.
The comprehensive analysis in Section 3 reveals that the zero-shot vision encoder $f_{vis}$ of CLIP is very sensitive to image corruption with increasing severity. Similarly, the performance of the text encoder $f_{txt}$ is invariant to the different text prompt templates and is also impractical to tune at test-time. Therefore, for effective adaptation, both the image and text features of CLIP need to be adapted to the incoming domain of test batches. We illustrate our proposed framework, BAT-CLIP, in Figure 3.
From unimodal adaptation to bimodal adaptation. The goal of TTA is to enhance a model’s performance in the current domain to ensure accurate predictions. To handle the complexities of CLIP adaptation to a specific domain of image corruption, we dissect our analysis of each encoder’s adaptation. Consider the unimodal adaptation of the vision encoder $f_{vis}$ via entropy minimization [46]. While the image features are adjusted for a specific test batch and domain, the text features remain fixed, still being optimized for the data from CLIP pre-training, as seen in Figure 3 (top row). Likewise, if only the text encoder $f_{txt}$ is updated in this manner, the generated text features may not align properly with the distribution of the incoming data, potentially causing misalignment between the image and text features, missing in the ideas of TPT [38] and VTE [9].
To benefit from the feature space and learn richer representations across both modalities, we propose adapting both $f_{vis}$ and $f_{txt}$ to a domain, enabling an input-aware knowledge transfer between the encoders and enhancing domain-specific adaptation. While the encoders can be updated based on entropy minimization, they come with inherent limitations. Though entropy models the prediction uncertainty, it does not guarantee increasing the likelihood of alignment between the image and text features. Additionally, due to the image corruption, for images within a test batch, there could be a possible overlap of visual features belonging to different classes [18]. The robustness of CLIP would be challenged in such a case. To address these limitations, we propose two loss components that facilitate a more effective bimodal adaptation of CLIP to new domains at test-time. The losses focus on maximizing alignment between the visual and text features while increasing separation between the visual features to learn good decision boundaries.
Projection matching between the visual and text features. We propose learning visual and text features that are domain-specific and mutually aware by jointly updating the encoders, as in Figure 3 (middle row), opposed to a unimodal update in TPT. From Eq. 1, the visual feature $v_{k,i}^{t}$ of the $k^{th}$ image in batch $x_{i}^{t}$ needs to have a high similarity with the text feature $z_{c}$ of class c for good alignment. To quantify this similarity, a possible direction is to project the visual feature onto the text feature i.e., compute the scalar projection — $v_{k,i}^{t}\cdot\hat{z}_{c}$, where $\hat{z}_{c}$ is the normalized text feature of $\hat{z}_{c}$ i.e., $\hat{z}_{c}=\frac{z_{c}}{||z_{c}||_{2}}$. Geometrically speaking, maximizing this projection leads to more similarity between the corresponding features and hence, a better alignment.
A possible approach would be to compute the projection for each class-specific visual feature in a test batch. However, within the batch, due to the image corruption, the pseudo-labels could be wrong/noisy. Instead, we propose modeling the projection of the class prototype with its corresponding text feature. A prototype is useful, in such a scenario, because it encompasses the entire class distribution without relying on individual visual features [39]. In particular, we compute a class prototype as,
$\bar{v}_{c}$ $=\frac{1}{\sum_{k=1}^{\mathcal{B}}\mathds{1}[\hat{y}=c]}\sum_{k=1}^{\mathcal{B}}\mathds{1}[\hat{y}=c]v_{k,i}^{t}$ (2)
[…]
i.e., we compute the mean feature of all the support visual features constituting a class c, based on the computed pseudo-labels (Eqs. 1 and 2), where $\hat{y}$ refers to the pseudo-labels computed via Eq. 1 and $\bar{v}_{c}$ is the class prototype of class c. Based on the class prototype, the projection matching loss is,
$\mathcal{L}_{pm}=\frac{1}{C}\sum_{c}\bar{v}_{c}\cdot\hat{z}_{c}$ (3)
Eq. 3 encourages a class prototype to have a larger projection on its corresponding text feature. In this way, during the adaptation of CLIP to a certain domain, both encoders learn to generate richer visual and text features with maximum alignment by maximizing Eq. 3. Figure 3 (middle row) shows such an illustration.
Inter-class separability between class prototypes. The projection matching loss introduced in Eq. 3 encourages the $f_{vis}$ and $f_{txt}$ encoders to produce domain-specific, well-aligned, and input-aware features via jointly updating the encoders. However, when TTA occurs at a batch level with image corruption, visual features within the batch could overlap, leading CLIP to poorly differentiate between classes. This ultimately hinders effective class separation. Hence, with a desire to obtain separation between visual features from different classes, as illustrated in Figure 3 (last row), we propose maximizing the distance between the class prototypes. Since the visual and text features align across modalities via Eq. 3, class separation, in addition, is needed for good generalization, robustness, and adaptation. So, we increase the cosine distance between the class prototypes and enhance the discriminative nature as,
$\mathcal{L}_{sp}=\sum_{l\in C}\sum_{c\in C}\mathds{1}[l\neq c](1-\text{cos}(\bar{v}_{c},\bar{v}_{l}))$ (4)
Optimization. TENT [46] optimizes the output logits by minimizing the entropy [37]. When applied to CLIP, the entropy, with output logits l, is defined as,
$\mathcal{L}_{ent}=-\sum_{c}p(l_{c})\log p(l_{c})$ (5)
where $p(l_{c})$ is the likelihood for class c that is computed via Eq. 1. The overall optimization objective for our approach is as,
$\underset{\phi_{v},\phi_{t}}{\arg\!\min}(\mathcal{L}_{ent}-\mathcal{L}_{pm}-\mathcal{L}_{sp})$ (6)
where $\phi_{v}$ and $\phi_{t}$ refer to the parameters of $f_{vis}$ and $f_{txt}$, respectively.
Motivation of choice of parameters. Through thorough experimentation, [50] demonstrates that fine-tuning LayerNorm parameters in attention-based models leads to yielding strong results on multimodal tasks. This also brings a reduction in trainable parameters. Updating normalization layers has also been deeply studied in tasks involving domain shifts [21]. Motivated by these works, we update the LayerNorm parameters of both CLIP encoders. This constitutes updating $\sim 0.044$% of all parameters. For every new task/corruption, we reset the model parameters following TENT [46] i.e., we perform a single domain TTA. We use a generic hand-crafted prompt template “a photo of a ¡CLS¿.” at all times, thus avoiding the need for a prompt search. In summary, our bimodal test-time adaptation approach jointly updates the LayerNorm parameters of the CLIP encoders that are optimized synergistically through loss components aware of the input domain, leading to a more robust multimodal learning process.
5 Experiments and Results
5.1 Online Test-Time Adaptation
Baselines. We compare our approach with zero-shot CLIP and other proposed online TTA methods using CLIP - TPT [38] and VTE [9]  [^2] We do not benchmark against WATT [29] since they update the model several times, for each prompt template, over a number of prompt templates.. For TPT, 63 random augmentations are generated for a single test image in the vision space. Based on a threshold, the marginal entropy of confident predictions is minimized to optimize the prompts. In the text space, VTE considers an ensemble of hand-crafted prompts for a test image, to get the final predictions (Eq. 1). Alongside this, they also consider multiple random augmentations of the test image, with frozen CLIP encoders. We also contrast our work against state-of-the-art TTA methods i.e., TENT [46], RoTTA [47], RPL [34], and SAR [28]. In essence, we adopt these approaches to use CLIP. Following the guidelines set by [9], we update only the vision encoder $f_{vis}$ for such a setup. The model predictions can then be computed as in Eq. 1. We mention the training details of each TTA method in the Supplementary.
Implementation Details. We query the text encoder $f_{txt}$ with a general and fixed hand-crafted prompt template “a photo of a ¡CLS¿.” for all the datasets, as motivated earlier. We optimize both the vision encoder $f_{vis}$ and text encoder $f_{txt}$. For CIFAR-10C, we use an AdamW optimizer at a learning rate of $10^{-3}$. Similarly, for CIFAR-100C and ImageNet-C, Adam and AdamW optimizers are respectively used, at a fixed learning rate of 5$\times$$10^{-4}$, with the model being reset after each task. The batch sizes are set to 200, 200, and 64 for the datasets, following various TTA benchmarks, at a corruption severity level of 5 for each task. As visual backbones, we report results on ViT-B/16 and ViT-B/32 [10], where all of our experiments are run on a single NVIDIA RTX A5000 GPU.
Method
Gaussian
Shot
Impulse
Defocus
Glass
Motion
Zoom
Snow
Frost
Fog
Brightness
Contrast
Elastic
Pixelate
JPEG
Mean
CIFAR-10C
ViT-B/16
37.92 41.7 54.42 71.75 40.89 67.93 73.62 73.89 77.35 70.22 84.45 62.36 53.81 47.65 59.43 61.16
TPT
37.74 42.24 60.57 72.88 44.80 69.69 75.37 75.96 78.84 72.12 85.68 62.04 58.90 55.14 62.64 63.64
VTE
42.42 46.26 64.23 71.10 45.58 68.50 73.66 76.75 78.27 71.02 85.28 57.24 59.54 60.59 61.85 64.15
Ours
61.13 64.09 65.76 80.51 54.96 80.65 81.94 83.04 84.19 80.84 88.95 82.15 69.16 62.68 66.64 73.85
Gain/Loss(%)
+18.71 +17.83 +1.53 +7.63 +9.38 +10.96 +6.57 +6.29 +5.35 +8.72 +3.27 +19.79 +9.62 +2.09 +4.00 +9.70
ViT-B/32
35.47 39.94 43.23 69.95 41.43 64.50 70.13 70.85 72.33 66.66 81.37 64.57 59.69 48.28 56.62 59.00
TPT
43.11 46.53 48.29 71.31 47.80 66.89 71.96 74.00 76.00 68.81 84.12 66.35 63.86 51.86 58.01 62.59
VTE
47.59 50.18 53.15 71.39 53.86 67.92 72.90 76.37 76.30 70.78 83.27 61.07 69.00 58.57 61.14 64.90
Ours
52.39 55.99 52.54 76.79 54.04 74.90 75.79 77.67 79.10 75.31 86.33 77.34 67.41 57.06 61.29 68.26
Gain/Loss(%)
+4.80 +5.81 -0.61 +5.40 +0.18 +6.98 +2.89 +1.30 +2.80 +4.53 +2.21 +10.99 -1.59 -1.51 +0.15 +3.36
CIFAR-100C
ViT-B/16
19.64 21.40 25.26 42.54 20.03 43.17 47.95 48.35 49.74 41.57 57.02 34.58 29.15 23.96 32.43 35.79
TPT
17.95 19.51 27.13 43.53 20.08 42.65 48.63 49.11 49.48 42.14 57.35 33.26 31.13 27.59 32.75 36.15
VTE
17.96 18.72 28.17 40.38 19.60 39.50 45.33 48.24 46.87 40.73 55.31 30.04 32.47 30.35 31.45 35.01
Ours
24.91 27.73 33.66 50.11 26.27 48.49 54.85 52.35 51.62 48.38 63.27 45.21 34.74 32.38 37.31 42.09
Gain/Loss(%)
+5.27 +6.33 +5.49 +6.58 +6.19 +5.32 +6.22 +3.24 +1.88 +6.24 +5.92 +10.63 +2.27 +2.03 +4.56 +5.94
ViT-B/32
16.23 17.83 17.57 39.07 17.63 38.55 43.81 42.32 43.46 39.71 50.32 29.34 28.74 22.85 29.42 31.79
TPT
16.08 17.65 17.54 39.21 19.47 38.91 44.01 43.45 44.46 40.15 50.93 27.77 30.91 23.36 29.55 32.23
VTE
16.84 18.33 18.94 39.63 22.88 39.13 43.80 44.56 44.88 39.21 49.37 28.37 34.13 26.87 30.12 33.14
Ours
21.35 24.71 22.32 46.26 23.07 44.64 50.12 47.23 46.88 44.92 58.55 38.52 34.56 27.73 33.19 37.60
Gain/Loss(%)
+4.51 +6.38 +3.38 +6.63 +0.19 +5.51 +6.11 +2.67 +2.00 +4.77 +7.62 +9.18 +0.43 +0.86 +3.07 +4.46
ImageNet-C
ViT-B/16
11.18 12.54 12.04 23.36 15.18 24.50 22.58 32.32 29.88 35.88 54.18 17.20 12.72 30.96 33.26 24.51
TPT
8.48 9.46 10.20 23.98 15.16 25.10 24.00 33.94 32.12 37.08 55.64 16.54 13.68 34.06 33.58 24.87
VTE
9.18 10.76 10.78 24.72 14.30 24.36 25.24 35.38 32.46 38.16 55.56 16.14 14.26 38.72 33.98 25.60
Ours
19.32 21.38 19.60 26.58 21.94 30.88 29.02 36.48 32.00 40.98 56.72 26.14 23.74 37.67 38.34 30.72
Gain/Loss(%)
+8.14 +8.84 +7.56 +1.86 +6.76 +5.78 +3.78 +1.10 -0.46 +2.82 +1.08 +8.94 +9.48 -1.05 +4.36 +5.12
ViT-B/32
12.88 13.04 12.90 24.42 11.86 22.72 20.20 25.70 25.84 30.28 50.54 17.32 18.96 32.20 29.12 23.20
TPT
12.04 12.64 12.52 25.38 12.28 22.68 20.78 26.36 26.64 30.78 51.02 16.50 19.90 33.62 30.62 23.58
VTE
11.96 12.32 13.44 25.06 11.70 22.58 22.40 27.38 27.02 32.28 51.52 16.84 19.94 34.80 32.82 24.14
Ours
16.84 18.20 16.10 25.04 20.90 28.90 25.24 29.42 27.18 36.02 50.18 17.66 27.68 36.20 35.42 27.39
Gain/Loss(%)
+3.96 +5.16 +2.66 -0.34 +8.62 +6.18 +2.84 +2.04 +0.16 +3.74 -1.34 +0.34 +7.74 +1.40 +2.60 +3.25
Table 2: Mean accuracy (%) on CIFAR-10C, CIFAR-100C, and ImageNet-C - TTA mean accuracy of the 15 corruptions (tasks) at a severity level of 5, using ViT-B/16 and ViT-B/32. + and - denote the absolute gain/loss w.r.t the next best performance. We contrast our results against zero-shot ViT-B/16 and -B/32, TPT [38], and VTE [9].
Gaussian Defocus Fog Pixelate
ViT-B/16 Refer to caption Refer to caption Refer to caption Refer to caption
Ours Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: BAT-CLIP yields more discriminative visual features that exhibit stronger alignment with their corresponding text features. The t-SNE [43] plots show visual ($\circ$) and text ($\bigstar$) features for CIFAR-10C, comparing zero-shot ViT-B/16 with our approach.
Results on CIFAR-10C, CIFAR-100C, and ImageNet-C. We present the TTA results in Table 2. Our method significantly improves the mean accuracy across all tasks and datasets for both visual backbones. Unlike TPT [38] and VTE [9], which process a single test image at a time and therefore incur substantial time, our approach operates at the batch level. This allows us to leverage the semantic relationships within the batch. By updating the parameters of $f_{vis}$ and $f_{txt}$ through the projection matching loss and enhancing the cosine distance between the class prototypes, our method effectively strengthens the alignment between visual and text features and learns discriminative visual features. In Figure 4, we show t-SNE [43] plots of visual and text features, of a few tasks, and compare them against zero-shot ViT-B/16 for CIFAR-10C. It highlights how our approach enhances alignment with text features, fosters better class separation, and forms more distinct clusters, leading to significant improvements compared to zero-shot CLIP. The Supplementary shows detailed t-SNE plots for CIFAR-10C and CIFAR-100C. Concerning compute time per task, on a difficult dataset like ImageNet-C, our method takes about 45 s, compared to 40 mins for TPT and 4 mins for VTE. This makes our approach deployment-friendly for quick classification with large improvements.
Dataset Backbone Zero-Shot TENT RoTTA RPL SAR Ours
CIFAR-10C ViT-B/16 61.16 62.06 61.89 61.52 67.37 73.85
ViT-B/32 59.00 56.35 59.57 56.51 65.13 68.26
CIFAR-100C ViT-B/16 35.79 37.96 36.12 38.47 41.19 42.09
ViT-B/32 31.79 31.61 31.84 30.00 37.81 37.60
ImageNet-C ViT-B/16 24.51 25.15 24.78 25.08 29.73 30.72
ViT-B/32 23.20 24.05 23.33 23.62 28.07 27.39
Table 3: Mean accuracy (%) of CLIP adopted for TTA approaches at a corruption severity level of 5, using ViT-B/16 and ViT-B/32.
Results against prior TTA methods. In Table 3, we compare our method with existing TTA approaches, all adopted for CLIP. We report mean accuracy across all tasks. While most TTA methods outperform zero-shot CLIP on average, our approach consistently matches or exceeds the performance of these methods across all datasets and visual backbones. Detailed per-task accuracies are provided in the Supplementary. With a ViT-B/16 backbone, TENT [46] increases the accuracy over zero-shot CLIP by 0.9%, 2.18%, and 0.64%, for the respective benchmark datasets. SAR [28] shows significant performance improvements across both vision backbones over prior TTA methods. Although it performs similarly to our method on most benchmarks, our approach outperforms SAR on CIFAR-10C, achieving improvements of 6.48% and 3.13% on the ViT-B/16 and ViT-B/32 backbones, respectively. While all the prior approaches operate solely in the vision space, our approach leverages CLIP’s joint vision-language feature space, resulting in superior TTA performance.
5.2 Ablation Study and Analysis
Refer to caption Refer to caption Refer to caption
a) CIFAR-10C b) CIFAR-100C c) ImageNet-C
Figure 5: Analysis on increasing # iterations: Mean accuracy for # iterations, on each test batch, using a ViT-B/16 backbone.
Adaptation for multiple iterations. We adapt CLIP at test-time with one update per batch. Here, we analyze adaptation for multiple iterations on a single batch. In Figure 5, we illustrate the mean accuracy, across all 15 tasks, using the ViT-B/16 backbone. Continuous adaptation of LayerNorm parameters, which normalize along the features of a single sample, can lead to overfitting as it fits to the noise patterns of the batch, ultimately degrading generalization to other domains. This could also lead to a decline in the loss of CLIP pre-trained knowledge. In the Supplementary, we report the post-adaptation results back on the source test sets - CIFAR10 and CIFAR100. Our approach still maintains a significant advantage over zero-shot CLIP evaluation. Despite a gradual decline in mean accuracy, the performance gap remains notably better, particularly on CIFAR-10C and ImageNet-C. This suggests that the proposed loss components enhance robustness, maintaining effectiveness under challenging conditions.
Backbone (ViT-B/16) CIFAR-10C CIFAR-100C ImageNet-C
$\mathcal{L}_{ent}$ 60.65 38.17 24.03
$\mathcal{L}_{pm}$ 61.23 36.63 22.83
$\mathcal{L}_{sp}$ 73.16 41.36 30.05
$\mathcal{L}_{ent}$ + $\mathcal{L}_{pm}$ 62.60 39.32 25.21
$\mathcal{L}_{ent}$ + 0.1 * ($\mathcal{L}_{pm}$ + $\mathcal{L}_{sp}$) 71.90 41.88 30.24
$\mathcal{L}_{ent}$ + 0.5 * ($\mathcal{L}_{pm}$ + $\mathcal{L}_{sp}$) 73.70 41.98 30.44
$\mathcal{L}_{ent}$ + $\mathcal{L}_{pm}$ + $\mathcal{L}_{sp}$ 73.85 42.09 30.72
Table 4: Ablation on loss components - Mean accuracy (in %).
Contribution of each loss component. In Table 4, we validate the effectiveness of the loss components by taking their different combinations. The mean accuracies across all the tasks are reported using ViT-B/16. The task-wise results are reported in the Supplementary. We observe that the addition of $\mathcal{L}_{pm}$ and $\mathcal{L}_{sp}$ largely improve model performance than entropy minimization via $\mathcal{L}_{ent}$. In fact, the addition of $\mathcal{L}_{sp}$ brings larger improvements, proving that $f_{vis}$ produces discriminative features. Interestingly, using just $\mathcal{L}_{ent}$ achieves better or comparable accuracy than zero-shot CLIP, as the encoders work in synergy while adapting to the domain.
Impact of bimodal adaptation. To show the effectiveness of bimodal adaptation, we ablate $\mathcal{L}_{pm}$ in Eq. 6, which is responsible for updating the text encoder $f_{txt}$. We report the mean classification results using the ViT-B/16 vision backbone in Table 5. As can be seen, we see a drop in accuracy when $f_{txt}$ is “frozen” i.e., when $\mathcal{L}_{pm}$ isn’t used, necessitating the need for bimodal adaptation of CLIP encoders.
Dataset $f_{vis}$ update $f_{txt}$ update Ours
CIFAR-10C 72.58
73.85
CIFAR-100C 41.00
42.09
ImageNet-C 29.88
30.72
Table 5: Ablation on $\mathcal{L}_{pm}$ to demonstrate the need of bimodal adaptation of CLIP encoders - using a ViT-B/16 backbone.
Refer to caption Refer to caption
Figure 6: Ablation on low batch size - CIFAR-10C and CIFAR-100C (left) and ImageNet-C (right) using ViT-B/16 backbone.
Low batch size setting. In Eq 3, the prototype of each class, via pseudo-labels of each test batch, can be affected by the batch size, with a lower size indicating the absence of a few classes. In Figure 6, we illustrate the mean accuracies on the benchmark datasets. As seen, BAT-CLIP still outperforms or performs comparably to TPT and VTE in low batch size settings. Since TTA happens at the batch-level of a specific domain, the prototypes are computed with the available pseudo-labels, aiding in the continual learning of the feature space as new batches come by.
Effect of different prompt templates. In Table 6, we show results with “relevant” prompt templates to show the independence of such prompt selection, at test-time. As seen, the performance gain over zero-shot ViT-B/32 is fairly large for all the prompt templates. Though TPT fine-tunes a pre-trained prompt on each image, and VTE uses an ensemble of prompts, our method is agnostic to the prompt template being used, making it favorable for real-time usage.
Prompt Template CIFAR-10C CIFAR-100C ImageNet-C
“a low contrast photo of a ¡CLS¿.”
68.53 (+7.81) 37.09 (+4.97) 27.31 (+3.71)
“a blurry photo of a ¡CLS¿.”
68.84 (+10.96) 36.80 (+5.33) 26.92 (+3.52)
“a photo of a big ¡CLS¿.”
67.49 (+10.10) 35.79 (+4.87) 25.64 (+3.29)
Table 6: Prompt template selection. + denotes the accuracy gain over zero-shot ViT-B/32.
6 Conclusion
In this paper, we propose BAT-CLIP, a bimodal online test-time adaptation framework for CLIP aimed at handling diverse image corruptions simulating real-time environments [15]. Our in-depth analysis of CLIP’s zero-shot performance under increasing corruption severity reveals significant shortcomings in generalization, highlighting the need for effective adaptation. While prior works on TTA for CLIP have predominantly been unimodal focusing on prompt-tuning [38] or prompt ensemble [9] with no model updates, our approach encourages learning rich class-separated visual features via vision encoder updates and strengthens alignment between the image class prototype and corresponding text feature via the text encoder updates. Empirical studies, including ablation results, demonstrate the robustness and substantial performance improvements of our approach over existing methods on benchmark corruption datasets — CIFAR-10C, CIFAR-100C, and ImageNet-C, through synergistic CLIP encoder updates.
Supplementary Material
In this work, we study the problem of test-time adaptation (TTA) of CLIP [31] towards common image corruptions and propose improved schemes for increasing the robustness of CLIP. We put forward a bimodal domain adaptation scheme, wherein we exploit the shared feature space of CLIP. In essence, leaning towards a more effective multimodal learning and adaptation method, we propose loss components that improve alignment between the class-specific visual prototype and corresponding text features via maximizing the projection. We also increase the cosine distance between the class prototypes to enhance discrimination between visual features. In this Supplementary, we provide additional insights and experimental results, that has been organized as follows,
  1. 1.
    Subsection 6.1 offers a detailed discussion of the datasets used, supplemented with visual illustrations.
  2. 2.
    To ensure full transparency, we outline the implementation details of all methods in section 6.2, including those for prior TTA approaches [46, 47, 34, 28] adapted for use with CLIP.
  3. 3.
    Section 6.3 presents further results and analysis:
    • In subsection 6.3.1, we explore the limitations of zero-shot CLIP when using ViT-B/32 and ViT-L/14 backbones under increasing image corruption severity.
    • Subsection 6.3.2 expands on task-specific accuracies, by adopting TTA approaches, for CIFAR-10C, CIFAR-100C, and ImageNet-C, along with a detailed performance discussion.
    • In subsections 6.3.3 and 6.3.4, we present a detailed loss ablation study and the post-adaptation zero-shot generalization on source test sets, respectively — essentially evaluating catastrophic forgetting of BAT-CLIP.
    • Lastly, we include task-wise t-SNE visualizations in subsection 6.3.5 for CIFAR-10C and CIFAR-100C, comparing our method against zero-shot CLIP (ViT-B/16), to illustrate the effectiveness of BAT-CLIP.
6.1 Datasets
We employ the CIFAR-10C, CIFAR-100C, and ImageNet-C datasets, for our experiments, as introduced by [15]. Each dataset includes 15 distinct types of image corruptions, referred to as tasks in a test-time adaptation setting, applied to the test sets of CIFAR10, CIFAR100 [17], and ImageNet [7]. These corruptions are applied at 5 different severity levels, ranging from mild to severe. For each task, CIFAR-10C and CIFAR-100C contain 10,000 test samples, whereas ImageNet-C has 5000 samples.
The image corruptions are categorized into four primary groups: noise, blur, weather, and digital distortions. Noise-based corruptions include Gaussian, Shot, and Impulse noise, which introduce random pixel-level variations. The blur category encompasses Defocus, Glass, Motion, and Zoom blur effects, all of which simulate different types of distorted imagery. Weather-related corruptions, such as Snow, Frost, and Fog, replicate environmental conditions that obscure image details. Lastly, digital distortions include effects like Brightness, Contrast, Elastic Transform, Pixelate, and JPEG compression, which reflect various forms of post-processing or compression artifacts that degrade image quality.
These corruption types, as proposed by [15], provide a comprehensive framework for assessing model robustness, which has been and is still being studied [16, 27, 30, 41, 24]. Their ability to emulate real-world image degradation scenarios is advantageous, allowing for a more realistic evaluation of a model’s robustness. We provide corruption visualizations, via an image example, in Figure 7. We urge the readers to check out [15] for further inspection.
Gaussian Shot Impulse Defocus Glass
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Motion Zoom Snow Frost Fog
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Brightness Contrast Elastic Pixelate JPEG
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 7: We provide visualizations of an image from ImageNet-C [15] for different corruption types, at an image severity level of 5.
Figure 8: Comparison of classification predictions across various methods (Zero-shot CLIP (ViT-B/16), VTE [9], TPT [38], and Ours) on ImageNet-C samples with Gaussian noise. Each row illustrates an example, displaying the ground truth (GT) label alongside the predictions from each method. Correct predictions are highlighted in green, while incorrect ones are marked in red. Our approach demonstrates enhanced robustness and higher accuracy, especially in challenging image corruption conditions.
Method
Gaussian
Shot
Impulse
Defocus
Glass
Motion
Zoom
Snow
Frost
Fog
Brightness
Contrast
Elastic
Pixelate
JPEG
Mean
CIFAR-10C
ViT-B/16
37.92 41.70 54.42 71.75 40.89 67.93 73.62 73.89 77.35 70.22 84.45 62.36 53.81 47.65 59.43 61.16
TENT
15.49 18.28 38.12 81.59 21.73 76.32 82.35 84.62 82.19 80.60 91.83 80.55 63.52 58.57 54.71 62.03
RoTTA
39.17 42.90 55.41 72.18 41.32 68.02 74.01 74.38 78.01 70.80 84.80 63.19 54.62 49.32 60.15 61.89
RPL
15.47 17.43 40.73 81.76 20.08 69.89 82.93 84.43 83.19 81.84 91.80 79.42 64.89 54.07 54.90 61.52
SAR
47.98 53.60 60.56 74.30 47.56 73.15 76.43 77.91 79.88 75.66 86.79 71.62 58.34 62.03 64.71 67.37
Ours
61.13 64.09 65.76 80.51 54.96 80.65 81.94 83.04 84.19 80.84 88.95 82.15 69.16 62.68 67.64 73.85
ViT-B/32
35.47 39.94 43.23 69.95 41.43 64.50 70.13 70.85 72.33 66.66 81.37 64.57 59.69 48.28 56.62 59.00
TENT
20.09 23.45 34.47 69.85 23.01 39.79 60.35 76.83 77.49 76.07 88.88 81.38 65.35 57.01 51.19 56.35
RoTTA
36.55 40.91 43.99 70.03 42.45 64.52 70.08 71.23 72.68 67.31 81.92 64.99 60.33 49.40 57.11 59.57
RPL
15.89 19.08 34.04 77.84 18.72 41.22 62.39 78.17 78.86 76.31 88.83 81.15 68.98 54.19 51.91 56.51
SAR
50.28 54.12 49.65 73.08 51.98 71.17 74.65 73.73 75.22 70.99 84.25 72.08 63.93 51.57 60.32 65.13
Ours
52.39 55.99 52.54 76.79 54.04 74.90 75.79 77.67 79.10 75.31 86.33 77.34 67.41 57.06 61.29 68.26
Table 7: Mean accuracy (%) on CIFAR-10C - TTA mean accuracy of the 15 corruptions (tasks) at a severity level of 5, using ViT-B/16 and ViT-B/32. We contrast our results against zero-shot backbones, TENT [46], RoTTA [47], RPL [34], and SAR [28].
Method
Gaussian
Shot
Impulse
Defocus
Glass
Motion
Zoom
Snow
Frost
Fog
Brightness
Contrast
Elastic
Pixelate
JPEG
Mean
CIFAR-100C
ViT-B/16
19.64 21.40 25.26 42.54 20.03 43.17 47.95 48.35 49.74 41.57 57.02 34.58 29.15 23.96 32.43 35.79
TENT
7.60 8.21 8.33 51.81 7.95 52.45 55.34 54.16 36.17 50.92 65.63 54.51 36.52 43.99 35.81 37.96
RoTTA
20.65 22.22 26.17 42.48 20.26 42.90 47.88 48.75 49.92 41.86 57.00 34.52 29.27 25.08 32.88 36.12
RPL
6.44 7.09 7.09 52.16 11.81 52.33 55.50 54.20 38.83 51.99 66.07 54.45 36.86 42.83 39.45 38.47
SAR
25.30 27.19 32.78 47.12 23.42 47.16 51.70 51.94 52.48 48.77 61.54 44.50 32.26 33.67 38.06 41.19
Ours
24.91 27.73 33.66 50.11 26.27 48.49 54.85 52.35 51.62 48.38 63.27 45.21 34.74 32.38 37.31 42.09
ViT-B/32
16.23 17.83 17.57 39.07 17.63 38.55 43.81 42.32 43.46 39.71 50.32 29.34 28.74 22.85 29.42 31.79
TENT
5.53 7.64 6.85 49.60 4.47 48.45 52.35 49.77 26.77 37.50 63.05 50.53 13.89 27.00 30.80 31.61
RoTTA
16.63 18.25 17.78 38.62 17.76 38.38 43.52 42.39 43.41 39.37 50.60 28.85 28.89 23.50 29.65 31.84
RPL
4.50 5.80 9.61 50.26 4.43 48.88 52.61 50.27 22.36 25.34 63.36 50.31 9.10 18.65 34.53 30.00
SAR
24.63 27.14 21.25 44.57 22.98 43.95 48.40 48.01 47.76 44.85 57.76 42.11 32.69 28.02 33.08 37.81
Ours
21.35 24.71 22.32 46.26 23.07 44.64 50.12 47.23 46.88 44.92 58.55 38.52 34.56 27.73 33.19 37.60
Table 8: Mean accuracy (%) on CIFAR-100C - TTA mean accuracy of the 15 corruptions (tasks) at a severity level of 5, using ViT-B/16 and ViT-B/32. We contrast our results against zero-shot backbones, TENT [46], RoTTA [47], RPL [34], and SAR [28].
Method
Gaussian
Shot
Impulse
Defocus
Glass
Motion
Zoom
Snow
Frost
Fog
Brightness
Contrast
Elastic
Pixelate
JPEG
Mean
ImageNet-C
ViT-B/16
11.18 12.54 12.04 23.36 15.18 24.50 22.58 32.32 29.88 35.88 54.08 17.20 12.72 30.96 33.26 24.51
TENT
5.14 5.70 7.44 25.22 19.34 26.80 24.16 33.56 30.42 37.74 54.24 22.50 13.90 35.02 36.08 25.15
RoTTA
11.34 12.96 12.32 23.38 15.50 24.66 22.90 32.56 30.02 35.98 54.32 17.20 12.80 31.06 33.46 24.78
RPL
9.04 10.04 10.96 24.40 17.40 26.28 23.76 32.70 30.62 36.64 54.04 19.38 13.24 33.14 34.60 25.08
SAR
17.96 20.46 20.68 25.72 23.04 29.52 26.04 34.92 32.74 39.00 55.00 27.14 19.64 36.66 37.50 29.73
Ours
19.32 21.38 19.60 26.58 21.94 30.88 29.02 36.48 32.00 40.98 56.72 26.14 23.74 37.68 38.34 30.72
ViT-B/32
12.88 13.04 12.90 24.42 11.86 22.72 20.20 25.70 25.84 30.28 50.54 17.32 18.96 32.20 29.12 23.20
TENT
9.18 8.50 10.42 26.02 15.72 26.06 21.64 27.12 26.18 31.60 50.58 22.28 20.12 34.06 31.30 24.05
RoTTA
13.10 13.30 13.02 24.48 11.96 22.86 20.28 26.06 26.06 30.24 50.46 17.34 19.16 32.44 29.18 23.33
RPL
11.68 10.98 12.10 25.68 13.24 23.98 20.84 26.32 26.12 30.86 50.62 19.30 19.48 33.14 29.92 23.62
SAR
19.82 20.36 20.92 25.78 20.40 28.34 23.10 28.12 28.38 34.74 51.10 24.60 24.38 36.54 34.40 28.07
Ours
16.84 18.20 16.10 25.04 20.90 28.90 25.24 29.42 27.18 36.02 50.18 17.66 27.58 36.20 35.42 27.39
Table 9: Mean accuracy (%) on ImageNet-C - TTA mean accuracy of the 15 corruptions (tasks) at a severity level of 5, using ViT-B/16 and ViT-B/32. We contrast our results against zero-shot backbones, TENT [46], RoTTA [47], RPL [34], and SAR [28].
6.2 Implementation Details
In this section, we summarize the implementation details of all the baseline methods that have been mentioned in the main paper, including ours. We build our approach on the open-source code base  [^3] https://github.com/mariodoebler/test-time-adaptation/tree/main. a standard TTA benchmark codebase, that also houses the hyperparameters and training details of all the prior TTA methods. CLIP-like models are used as provided by OpenCLIP. Only the vision encoder is updated for online TTA methods [46, 47, 34, 28] adopted for CLIP.
BAT-CLIP (Ours): For domain-specific test adaptation, we conducted experiments using ViT-B/16 and ViT-B/32 [10] as the vision backbones. For CIFAR-10C, both the vision encoder ($f_{vis}$) and text encoder ($f_{txt}$) were updated using the AdamW optimizer with a learning rate of 10-3. Similarly, for CIFAR-100C and ImageNet-C, we employed the Adam optimizer and AdamW optimizer, respectively, with a learning rate of 5$\times$10-4. The batch size $\mathcal{B}$ used was set to 200 for CIFAR-10C and CIFAR-100C, and 64 for ImageNet-C. Throughout, the prompt template is fixed to “a photo of a ¡CLS¿.”.
TPT [38]: For each test image, 63 augmentations are generated based on random resized crops, yielding a batch of 64 images, in addition to the original test image. The prompt/context vectors are initialized based on “a photo of a ¡CLS¿.” and tokenized using pre-trained CLIP weights. The confidence threshold is set to 10% i.e., the marginal entropy over the 10% confident samples is minimized. For all the datasets, we follow their core implementation and optimize the prompt vectors using an AdamW optimizer with a learning rate of 5$\times$10-3.
VTE [9]: In VTE, an ensemble of different prompt templates is considered based on the idea of [31]. An example of templates includes “a photo of a ¡CLS¿.”, “a sketch of a ¡CLS¿.”, “a painting of a ¡CLS¿.”, etc. The prompt templates are then averaged. On the vision side, similar to TPT [38], a batch of random augmentation is created for a test image with no model updates.
TENT [46]: We follow all the hyperparameters that TENT provides in their official implementation  [^4] https://github.com/DequanWang/tent. To update the vision encoder, we use Adam as the optimizer with a learning rate of 10-3 for CIFAR-10C and CIFAR-100C. For ImageNet-C, we update using SGD with a learning rate of 25$\times$10-5.
RoTTA [47]: For fairness, the batch sizes are set to 200 for the CIFAR datasets and 64 for ImageNet-C. The vision encoder is updated based on the Adam rule with a learning rate of 10-3. The capacity of the memory bank is set to 64, for all the datasets. Following the notations in the paper, $\alpha$ = 0.05, $\delta$ = 0.1, $\nu$ = 0.001, $\lambda_{t}$ and $\lambda_{u}$ = 1.0. We implement the details exactly as described in their main paper.
RPL [34]: We use an Adam optimizer with a learning rate of 10-3 for CIFAR-10C and CIFAR-100C. For ImageNet-C, the update rule is SGD with a learning rate of 5$\times$10-4. To compute the generalized cross-entropy loss, q is set to 0.8 for all the datasets.
SAR [28]: The training details/hyperparameters for SAR are the same as RPL [34] for CIFAR-10 and CIFAR-100. For ImageNet-C, the learning rate is set to 25$\times$10-5 with an SGD update rule. The entropy threshold $E_{0}$ is 0.4xln(C), where C is the number of classes. $\rho$ is set to a default of 0.05. The moving average factor is 0.9 for $e_{m}$ and $e_{0}$ is set to 0.2. All parameters are the same as in [28].
Figure 9: Task-wise mean accuracy (%) of zero-shot CLIP across different corruption severity levels. [Top]: ViT-B/32 backbone. [Bottom]: ViT-L/14 backbone. The dashed lines indicate the performance of zero-shot CLIP (w/ respective visual backbones) on the corresponding source datasets.
6.3 Additional Results
In the following subsections, we provide additional results and discussions.
6.3.1 Zero-shot performance analysis of ViT-B/32 and ViT-L/14
In the main paper, we analyze and evaluate the zero-shot performance of ResNet-101 (RN101) [14] and ViT-B-16 [10] and conclude that such CLIP backbones are extremely sensitive, in terms of classification accuracy, to increasing severity levels of image corruption. This could be a major concern in situations involving real-time deployment of CLIP. Here, we present a similar analysis in Figure 9, using ViT-B/32 and ViT-L/14 as backbones. Our analysis, from the main paper, carries forward. To summarise, irrespective of the CLIP visual backbone, the robustness towards image corruption is limited. The classification performance degrades with an increase in the severity of corruption in an image.
Method
Gaussian
Shot
Impulse
Defocus
Glass
Motion
Zoom
Snow
Frost
Fog
Brightness
Contrast
Elastic
Pixelate
JPEG
Zero-Shot
ViT-B/16
90.1
Ours
84.51 84.29 88.69 88.48 86.44 86.46 87.04 91.38 91.01 90.22 90.87 88.12 88.13 74.74 87.48 87.19 (mean)
ViT-B/32
88.3
Ours
67.41 68.28 84.23 80.50 75.37 79.75 78.55 87.67 86.36 85.83 90.04 80.89 81.56 82.74 82.36 80.77 (mean)
Table 10: Zero-shot performance on CIFAR10 (source) after adaptation of BAT-CLIP on a task.
Method
Gaussian
Shot
Impulse
Defocus
Glass
Motion
Zoom
Snow
Frost
Fog
Brightness
Contrast
Elastic
Pixelate
JPEG
Zero-Shot
ViT-B/16
66.6
Ours
67.09 67.12 67.08 70.31 63.70 66.83 70.12 70.10 68.19 69.55 71.05 66.49 65.13 59.98 67.60 67.36 (mean)
ViT-B/32
62.3
Ours
45.99 46.90 60.86 63.62 57.73 59.59 61.66 66.21 61.50 63.61 66.39 57.05 62.13 62.82 65.37 60.09 (mean)
Table 11: Zero-shot performance on CIFAR100 (source) after adaptation of BAT-CLIP on a task.
6.3.2 Results against prior TTA methods
We adopt prior TTA methods [46, 47, 34, 28] for CLIP and present the fine-grained results i.e., the task-wise mean accuracy to demonstrate the efficacy of our method (the higher the better). The results are reported in Tables 7, 8, and 9. Across all the backbones and datasets, we see that our method BAT-CLIP, achieves the best or comparable performance against all the baseline TTA approaches adopted for CLIP.
6.3.3 Detailed results from the loss ablation study
In the main paper, we provide ablation of loss components and their combinations i.e., the mean accuracy across all the tasks for ViT-B/16 and ViT-B/32 on the benchmarks corruption datasets. Here, we provide additional task-wise accuracy in Table 12. The addition of loss components $\mathcal{L}_{pm}$ and $\mathcal{L}_{sp}$ help CLIP in adapting its feature space to a specific domain/corruption.
Method
Gaussian
Shot
Impulse
Defocus
Glass
Motion
Zoom
Snow
Frost
Fog
Brightness
Contrast
Elastic
Pixelate
JPEG
Mean
CIFAR-10C
ViT-B/16
$\mathcal{L}_{ent}$
14.62 17.29 49.25 81.06 20.23 74.27 81.10 84.25 81.93 80.93 91.86 78.92 53.09 51.18 49.79 60.65
$\mathcal{L}_{pm}$
41.82 45.66 52.88 70.13 39.29 66.61 71.27 71.49 74.71 69.20 80.98 72.89 55.58 55.55 50.44 61.23
$\mathcal{L}_{sp}$
62.47 65.43 63.41 79.96 52.73 80.02 81.38 82.35 83.44 80.46 88.85 81.22 67.77 60.52 67.74 73.16
$\mathcal{L}_{ent}$+$\mathcal{L}_{pm}$
16.58 19.89 42.69 79.45 23.41 77.03 80.95 81.74 78.45 80.66 90.52 82.55 62.56 64.35 58.16 62.60
$\mathcal{L}_{ent}$ + 0.1 *($\mathcal{L}_{pm}$ + $\mathcal{L}_{sp}$)
44.92 51.41 63.30 80.65 50.79 79.83 83.13 83.59 83.89 81.91 89.56 83.16 67.78 66.69 67.87 71.90
$\mathcal{L}_{ent}$ + 0.5 *($\mathcal{L}_{pm}$ + $\mathcal{L}_{sp}$)
58.81 63.85 65.99 80.26 53.90 80.30 82.30 83.08 84.04 81.66 89.19 82.67 68.29 62.70 68.52 73.70
$\mathcal{L}_{ent}$+$\mathcal{L}_{pm}$+$\mathcal{L}_{sp}$
61.13 64.09 65.76 80.51 54.96 80.65 81.94 83.04 84.19 80.84 88.95 82.15 69.16 62.68 67.64 73.85
CIFAR-100C
ViT-B/16
$\mathcal{L}_{ent}$
7.71 10.05 11.52 49.42 12.49 49.36 53.79 54.11 50.76 49.92 64.32 47.07 33.40 38.63 39.95 38.17
$\mathcal{L}_{pm}$
19.88 24.06 21.26 45.57 22.66 43.66 49.37 46.05 45.96 43.69 57.76 37.33 33.66 25.85 32.65 36.63
$\mathcal{L}_{sp}$
24.69 27.28 33.62 49.08 25.29 47.84 53.86 51.70 50.93 46.93 62.57 44.76 33.88 31.89 36.05 41.36
$\mathcal{L}_{ent}$+$\mathcal{L}_{pm}$
12.26 12.62 13.14 48.90 26.22 48.99 53.10 53.10 52.43 49.44 63.36 46.78 33.27 37.77 38.36 39.32
$\mathcal{L}_{ent}$ + 0.1 *($\mathcal{L}_{pm}$ + $\mathcal{L}_{sp}$)
24.99 27.10 32.95 49.92 25.91 48.42 54.43 52.87 51.57 47.70 63.68 45.25 39.74 31.58 37.16 41.88
$\mathcal{L}_{ent}$ + 0.5 *($\mathcal{L}_{pm}$ + $\mathcal{L}_{sp}$)
25.24 27.59 33.41 50.05 25.73 48.55 54.44 52.85 51.80 47.81 63.75 45.08 34.63 31.67 37.17 41.98
$\mathcal{L}_{ent}$+$\mathcal{L}_{pm}$+$\mathcal{L}_{sp}$
24.91 27.73 33.66 50.11 26.27 48.49 54.85 52.35 51.62 48.38 63.27 45.21 34.74 32.38 37.31 42.09
ImageNet-C
ViT-B/16
$\mathcal{L}_{ent}$
0.90 1.06 1.16 29.12 13.02 32.14 27.34 35.32 11.14 40.92 56.90 23.78 7.78 39.62 40.22 24.03
$\mathcal{L}_{pm}$
10.20 11.40 10.74 19.56 15.18 20.06 19.28 27.66 29.72 34.10 53.50 22.66 13.80 24.38 30.26 22.83
$\mathcal{L}_{sp}$
19.32 20.98 19.26 25.90 21.22 30.06 28.56 35.22 31.34 40.36 55.20 25.64 23.68 36.90 37.18 30.05
$\mathcal{L}_{ent}$+$\mathcal{L}_{pm}$
0.90 1.16 1.30 28.90 17.04 31.56 26.24 36.26 12.22 42.12 57.92 30.34 10.36 40.66 41.20 25.21
$\mathcal{L}_{ent}$ + 0.1 *($\mathcal{L}_{pm}$ + $\mathcal{L}_{sp}$)
19.48 21.14 19.16 26.80 20.70 29.80 29.32 35.92 30.68 41.04 55.80 25.66 22.50 37.68 37.92 30.25
$\mathcal{L}_{ent}$ + 0.5 *($\mathcal{L}_{pm}$ + $\mathcal{L}_{sp}$)
19.56 21.38 19.16 26.96 21.26 30.06 29.24 35.92 31.26 41.34 56.32 25.74 22.58 37.94 37.92 30.44
$\mathcal{L}_{ent}$+$\mathcal{L}_{pm}$+$\mathcal{L}_{sp}$
19.32 21.38 19.60 26.58 21.94 30.88 29.02 36.48 32.00 40.98 56.72 26.14 23.74 37.68 38.34 30.72
Table 12: Task-wise loss ablation results (accuracy) on CIFAR-10C, CIFAR-100C, and ImageNet-C.
6.3.4 Post-adaptation results on source test sets
Thanks to the natural language supervision and also due to the pre-training on large amounts of (image, text) pairs, CLIP has shown strong generalization capabilities. However, for an efficient adaptation to a downstream task, fine-tuning the full model is infeasible due to large model updates. The primary reason is the loss of useful pre-trained knowledge of CLIP, which could eventually lead to overfitting to a downstream task. However, for attention-based models, tuned for multimodal tasks, [50] show that tuning the LayerNorm parameters leads to strong results. Inspired by [50], in our bimodal test-adaptation scheme, we update the LayerNorm parameters of both CLIP encoders, to a specific corruption task, which makes it parametric-efficient. We perform a single-domain TTA or adapt CLIP, at test-time, to a single domain only and then reset the parameters. Now, with continual adaptation to a certain corruption task, it gets difficult to preserve CLIP’s pre-trained knowledge since the normalization parameters begin to overfit to this domain. Then, a natural question arises -
Given that CLIP has been adapted to a specific corruption task, will the zero-shot generalization still hold back on its source test set?
In this crucial experiment, we challenge our BAT-CLIP, and evaluate its zero-shot generalization performance back on the source test set, to check the preservation of pre-trained CLIP Knowledge. After the adaptation of CLIP on each corruption task, we report the adapted model’s zero-shot performance on its corresponding source test set. We report results for CIFAR-10C and CIFAR-100C in Tables 10 and 11, using ViT-B/16 and ViT-B/32 backbones. For all of the results, we use the prompt template “a photo of a ¡CLS¿.”. As an example, for CIFAR-10C, upon adaptation of CLIP to Gaussian noise following our approach, we report the adapted model’s zero-shot accuracy on its source test set - CIFAR10 test set.
From Table 10, we observe that, on average, there is a 2.91% drop in accuracy compared to a zero-shot evaluation using pre-trained CLIP ViT-B/16. Similarly, for ViT-B/32, we see a drop of about 7.53% in mean accuracy. In Table 11, for CIFAR-100C using a ViT-B/16 backbone, we see an improvement of 0.76% in mean accuracy.
On the whole, we conclude that since the adaptation for a task happens over multiple test batches, the zero-shot performance back on the source data largely depends on the distribution of the image corruption. Overall, ViT-B/16 visual backbones preserve larger amounts of CLIP pre-trained knowledge. This proves the effectiveness of our method BAT-CLIP, on average.
6.3.5 t-SNE visualizations of CIFAR-10C and CIFAR-100C
In this section, we provide illustrations of task-wise t-SNE [43] plots for CIFAR-10C and CIFAR-100C and compare them against zero-shot ViT-B/16. The results are in Figures 10, 11, 12, and 13. Across all corruptions/tasks, BAT-CLIP learns strong discriminative visual features with a strong image-text alignment and class-level separation. ImageNet-C has 1000 classes, so, we do not provide t-SNE plots to avoid complications. However, the analysis and results carry forward.
Gaussian Shot Impulse Defocus Glass
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Motion Zoom Snow Frost Fog
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Brightness Contrast Elastic Pixelate JPEG
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 10: BAT-CLIP (w/ ViT-B/16): The t-SNE plots show visual ($\circ$) and text ($\bigstar$) features for CIFAR-10C.
Gaussian Shot Impulse Defocus Glass
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Motion Zoom Snow Frost Fog
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Brightness Contrast Elastic Pixelate JPEG
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 11: Zero-shot ViT-B/16: The t-SNE plots show visual ($\circ$) and text ($\bigstar$) features for CIFAR-10C.
Gaussian Shot Impulse Defocus Glass
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Motion Zoom Snow Frost Fog
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Brightness Contrast Elastic Pixelate JPEG
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 12: BAT-CLIP (w/ ViT-B/16): The t-SNE plots show visual ($\circ$) and text ($\bigstar$) features for CIFAR-100C.
Gaussian Shot Impulse Defocus Glass
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Motion Zoom Snow Frost Fog
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Brightness Contrast Elastic Pixelate JPEG
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 13: Zero-shot ViT-B/16: The t-SNE plots show visual ($\circ$) and text ($\bigstar$) features for CIFAR-100C.

### v2

@@
-Enhancing Robustness of CLIP to Common Corruptions through Bimodal Test-Time Adaptation
+BATCLIP: Bimodal Online Test-Time Adaptation for CLIP
 Abstract
 Although open-vocabulary classification models like Contrastive Language Image Pretraining (CLIP) have demonstrated strong zero-shot learning capabilities, their robustness to common image corruptions remains poorly understood.
-Through extensive experiments, we show that zero-shot CLIP lacks robustness to common image corruptions at increasing severity levels during test-time, necessitating the adaptation of CLIP to unlabeled corrupted images using test-time adaptation (TTA).
+Through extensive experiments, we show that zero-shot CLIP lacks robustness to common image corruptions during test-time, necessitating the adaptation of CLIP to unlabeled corrupted images using test-time adaptation (TTA).
 However, we found that existing TTA methods have severe limitations in adapting CLIP due to their unimodal nature.
-To address these limitations, we propose BAT-CLIP, a bimodal TTA method specially designed to improve CLIP’s robustness to common image corruptions.
-The key insight of our approach is not only to adapt the visual encoders for better image feature extraction but also to strengthen the alignment between image and text features by promoting a stronger association between the image class prototype, computed using pseudo-labels, and the corresponding text feature.
-We evaluate our approach on benchmark image corruption datasets and achieve state-of-the-art results in TTA for CLIP, specifically for domains involving image corruption.
-Particularly, with a ViT-B/16 vision backbone, we obtain mean accuracy improvements of 9.7%, 5.94%, and 5.12% for CIFAR-10C, CIFAR-100C, and ImageNet-C, respectively.
+To address these limitations, we propose BATCLIP, a bimodal online TTA method designed to improve CLIP’s robustness to common image corruptions.
+The key insight of our approach is not only to adapt the visual encoders for improving image features but also to strengthen the alignment between image and text features by promoting a stronger association between the image class prototype, computed using pseudo-labels, and the corresponding text feature.
+We evaluate our approach on benchmark image corruption datasets and achieve state-of-the-art results in online TTA for CLIP.
+Furthermore, we evaluate our proposed TTA approach on various domain generalization datasets to demonstrate its generalization capabilities.
 1 Introduction
@@
 (b) VTE [10]
-(c) BAT-CLIP (Ours)
+(c) BATCLIP (Ours)
 (d) Performance comparison
-Figure 1: Comparison of BAT-CLIP with other online TTA approaches using CLIP: TPT [38], and VTE [9].
-a) TPT optimizes text prompts only for a single test image, making it unimodal.
-b) VTE considers an ensemble of prompts without model updates.
+Figure 1: Comparison of BATCLIP with other online TTA approaches using CLIP.
+a) TPT [45] optimizes text prompts only for a single test image, making it unimodal.
+b) VTE [10] considers an ensemble of prompts without model updates.
 Both methods consider the generation of multiple augmentations of the test image.
-c) BAT-CLIP is a bimodal approach, that adapts the LayerNorm parameters of the vision and text encoders, maximizing alignment between class prototypes and text features while increasing the inter-class separability of prototypes.
+c) BATCLIP is a bimodal approach, that adapts the LayerNorm parameters of the vision and text encoders, maximizing alignment between class prototypes and text features while increasing the inter-class separability of prototypes.
 The emergence of large pre-trained vision-language models (VLMs), such as CLIP [38], has led to their widespread adoption in various visual recognition tasks, including segmentation [24, 29], detection [2, 27], classification [61, 60], and image generation [52, 39, 40].
@@
 For example, models deployed for safety-critical applications like autonomous driving [1], could face rapid distributional shifts of blurriness, pixel changes, snowy nights, or other weather conditions [42].
-For instance, our findings on the zero-shot performance of CLIP with a ResNet-101 [14] vision backbone reveals that the accuracy on the test set of CIFAR100 [17] with Gaussian noise of severity level 5, plummets to 10.79% from 49% on the clean set.
+In particular, our findings on the zero-shot performance of CLIP with a ResNet-101 [17] vision backbone reveals that the accuracy on the test set of CIFAR100 [22] with Gaussian noise of severity level 5, plummets to 10.79% from 49% on the clean set.
 Similar trends are observed with ViT-B/16, -B/32, and -L/14 [11] as backbones.
-Such behaviors can result in significant performance degradation of models when encountering image corruption in real-world scenarios, particularly in applications like self-driving cars.
-This challenge is not unique to CLIP; it reflects a broader issue in computer vision tasks [7, 11, 12, 5], which often rely on the assumption that training and test data share the same distribution.
-When test distributions differ, model adaptation becomes essential to accommodate this shift and maintain accurate predictions.
-In response to this issue, test-time adaptation (TTA) has garnered significant interest [46, 8].
-TTA aims to adapt a pre-trained model to unlabeled batches of test data from varying domains in an online fashion—updating the model for each test batch without access to the source dataset.
-Several TTA methods proposed in the literature have been effective in mitigating domain shifts [46, 36, 34, 28, 42, 3].
+The performance degradation caused by image corruption can have significant consequences in real-world scenarios, particularly in safety-critical applications like self-driving cars.
+One effective approach to improving model performance under distribution shift is test-time adaptation (TTA) [53, 9], where a pre-trained model is directly adapted to unlabeled test batches, without access to the source dataset.
+Moreover, given the abundance of unlabeled data available in the wild, there is a growing need for online adaptation to streaming data, with a single forward pass, ensuring privacy, real-time performance [31], and preventing catastrophic forgetting [13] of source knowledge.
+Several online TTA methods proposed in the literature have been effective in mitigating domain shifts [53, 43, 41, 35, 48, 4].
 In this paper, we use the terms distributions and domains interchangeably.
-While there have been few new works on using CLIP for online TTA [38, 9], they come with certain limitations.
+While there have been few works on using CLIP for online TTA [45, 10, 21], they come with certain limitations.
 For example, test-time prompt tuning (TPT) [45] tunes the text prompts on the text encoder alone and generates multiple random augmented views, for each test image.
@@
 Another recent approach, Vision-Text-Space Ensemble (VTE) [10], uses an ensemble of different prompts as input to CLIP’s text encoder while keeping both the encoders frozen.
-Since the vision encoder is frozen, it struggles to adapt images with severe noise effectively.
+Since the vision encoder is frozen, it struggles to effectively adapt to images with severe noise.
 While TPT effectively improves CLIP’s test generalization by dynamically tuning the text prompts, it remains primarily a unimodal approach.
 This limits the capacity of a multimodal model like CLIP to fully leverage its multimodal nature for adaptation.
-Specifically, it prevents the encoders from jointly adjusting their features, resulting in suboptimal alignment between the visual and text modalities.
-During TTA, there is no transfer of knowledge, via gradients, between the visual and text encoders, as the adaptation process focuses solely on one modality.
+Specifically, it prevents the encoders from jointly adjusting their features, resulting in suboptimal alignment between the visual and text modalities after adaptation.
 For instance, when a test image includes common corruptions, the text prompts adapt, but the vision encoder features remain fixed.
 As a result, the learned prompts lack awareness of the test image distribution, leading to a less effective adaptation.
-To address these core limitations, we propose, BAT-CLIP, a bimodal adaptation approach of CLIP for online TTA, where both the visual and text encoders are jointly adapted by exploiting CLIP’s shared feature space of images and text.
+To address these core limitations, we propose, BATCLIP, a bimodal adaptation approach of CLIP for online TTA specifically, where both the visual and text encoders are jointly adapted by exploiting CLIP’s shared feature space of images and text.
 The overall objective is to achieve a strong alignment between image features and text features to enable more effective multimodal learning and adaptation.
-The adaptation procedure is two-fold: 1) Inspired by [47] for efficient fine-tuning, we adapt the norm layers i.e., LayerNorm parameters of both encoders.
+The adaptation procedure is two-fold: 1) Inspired by [54, 58] for efficient fine-tuning, we only adapt the norm layers, i.e., LayerNorm parameters of both encoders.
 However, such a model update does not consider the alignment of the encoder features.
-So, to improve the alignment between class-specific visual and text features, we introduce a projection matching loss that maximizes the projection of the visual class prototypes with their corresponding text features.
-2) To learn more discriminative visual features, we increase the cosine distance between the class prototypes, promoting a more distinct separation in the image feature space.
-Furthermore, our method is general-purpose, as it does not rely on hand-crafted prompt templates or their ensembles, unlike VTE.
-We leverage batches of test samples for TTA, rather than focusing on single-image adaptation, which makes our method fast.
+Therefore, to improve the alignment between class-specific visual and text features, we introduce a projection matching loss that maximizes the projection of the visual class prototypes with their corresponding text features.
+2) To learn more discriminative visual features, we enhance the cosine distance between class prototypes, computed using pseudo-labels, to promote a more distinct separation in the image feature space.
+Our method is general-purpose, with no reliance on multiple prompt templates [36, 15, 34] or their ensembles, unlike VTE.
+We leverage batches of test samples for TTA, rather than focusing on single-image test adaptation, which makes our method fast.
 In the Supplementary, we show several examples of classification results, as a comparison.
-Our novelty lies in performing a bimodal online test-time adaptation (TTA) of CLIP for classification tasks, to improve CLIP’s robustness.
-We draw a comparison of our method against TPT and VTE in Figure 1.
+Figure 1(d) shows that with a bimodal online test-time adaptation (TTA) setup for CLIP, the proposed approach achieves a significant improvement compared to TPT and VTE.
 Our main contributions are as follows:
-  • We start with an in-depth analysis of CLIP’s zero-shot performance across different visual backbones on common image corruptions, at test-time, for varying severity levels.
-While CLIP shows strong performance on clean images, its performance drops notably on corrupted images.
-  • To address the unimodal limitations highlighted earlier, our proposed bimodal adaptation of CLIP encoders at test-time aims to enhance alignment by maximizing the projection of class-wise prototypes onto their corresponding text features.
-Simultaneously, we increase the cosine distance between class prototypes to encourage the learning of more discriminative features, making the test adaptation process more flexible and robust.
-  • We conduct extensive experiments benchmarking our method against prior TTA approaches using CLIP, i.e., TPT and VTE, for online test-time adaptation.
-We also compare BAT-CLIP against four TTA baselines, adopted for CLIP.
-Our approach achieves state-of-the-art performance for CLIP adaptation to common corruptions.
+  • We start with an in-depth analysis of CLIP’s zero-shot performance across different visual backbones on an established benchmark on common image corruptions [18], at test-time, for varying severity levels.
+We found that while CLIP shows strong performance on clean images, its performance drops notably on corrupted images.
+  • To address the unimodal limitations highlighted earlier, we propose BATCLIP, a bimodal online test-time adaptation method for CLIP encoders, designed to enhance alignment by maximizing the projection of class-wise prototypes onto their corresponding text features.
+Simultaneously, we increase the cosine distance between class prototypes to encourage learning of more discriminative features, making the test adaptation process more flexible and robust.
+  • We conduct extensive experiments benchmarking the proposed method against established TTA baselines and others using CLIP on CIFAR-10C, CIFAR-100C, and ImageNet-C [18].
+To generalize our method beyond common corruptions, we evaluate on various domain generalization datasets [14]-OfficeHome [51], PACS [25], VLCS [12], and Terra Incognita [3].
+Overall, our approach achieves state-of-the-art performance for CLIP adaptation at test-time on various domain shifts.
 2 Related Works
 Online Test-Time Adaptation (TTA).
-The objective of TTA is to adapt a pre-trained model, with no access to the source data, to incoming batches of unlabelled test data of a specific domain [46, 42, 36, 34, 28, 3, 49, 6].
+The objective of online TTA is to adapt a pre-trained model, with no access to the source data, to incoming batches of unlabelled test data of a specific domain [53, 48, 43, 41, 35, 4, 57, 6].
 In this approach, the model is reset to its pre-trained state after adapting to each target domain.
@@
 However, none of these works utilize CLIP for TTA.
-TTA using CLIP.
-Lately, CLIP has been finding extensive applications for TTA.
-TPT [38] was the first work to propose prompt tuning using CLIP at test-time in an online manner.
-However, this method is computationally intensive for generating multiple views, per image.
-Additionally, a key limitation of this approach is its reliance on using CLIP’s default hand-crafted template for prompt initialization, with resets after adaptation to each image.
-A similar line of work was done in VTE [9] where ensembles are created in the text and vision space without any CLIP parameter update.
-[40] focus on single-image TTA for out-of-distribution detection and test generalization [22, 19].
-WATT [29] borrows the idea of “test-time training” i.e., over several optimization steps, it adapts CLIP’s vision encoder using multiple text prompt templates, for a batch of test images.
-However, this TTA setting differs from ours where we update the model only once on a single test batch.
-SwapPrompt [26] explores the idea of self-supervised contrastive learning for prompt learning.
-HisTPT [48] relies on a memory bank, to store features of recent test samples, for prompt learning.
-Such a setup is impractical in real-time and is expensive.
-To summarize — while TPT is unimodal, our approach is bimodal, involving on-the-fly joint optimization of both encoders for test adaptation, leading to a strong multimodal alignment between features.
-We use a generic prompt template, avoiding prompt engineering and enabling real-time deployment.
+[…]
+Another set of TTA works using CLIP [36, 15, 34] leverage information from multiple prompt templates.
+WATT [36] adapts CLIP’s vision encoder, weight-averaging the adapted weights from multiple prompt templates, for each test batch, over several optimization steps.
+Similarly, CLIPArTT [15] enhances text supervision from the support of pseudo-labels.
+Mishra et al.
+[34] pose pseudo-labeling as an Optimal Transport [7] problem by distilling knowledge from multiple text prototypes, for several iterations.
+Please note that the TTA setups here largely differ from our online TTA setup [53].
+While promising, these approaches have key limitations.
+Firstly, multiple gradient descent steps on a test batch can result in large parameter shifts, leading to the loss of CLIP’s source knowledge i.e., catastrophic forgetting [13].
+No regularization strategies have been proposed to restrict the loss of CLIP’s pretrained knowledge.
+Secondly, these methods are not favorable for real-time deployment where fast and on-the-fly test-adaptation must meet privacy and memory constraints.
+Thirdly, relying on multiple prompt templates, at test-time, is impractical.
+We discuss this in Section 3.
 3 Zero-shot performance analysis of CLIP to common image corruptions
@@
 Vision Backbones.
-We evaluate the robustness with a ResNet-101 (RN101) vision backbone [14] and three Vision Transformer backbones (ViT-B/16, ViT-B/32, ViT-L/14) [10].
-The models are tested on their zero-shot classification performance i.e., accuracy.
+We evaluate the robustness with a ResNet-101 (RN101) vision backbone [17] and three ViT backbones (ViT-B/16, ViT-B/32, and ViT-L/14) [11].
+The models are tested on their zero-shot classification performance.
 Datasets.
-For all the experiments in this paper, we utilize the CIFAR-10C, CIFAR-100C, and ImageNet-C [15] datasets, which are standard benchmark datasets to evaluate the robustness of classification models.
+For all the experiments in this paper, we utilize the CIFAR-10C, CIFAR-100C, and ImageNet-C [18] datasets, which are standard benchmark datasets to evaluate the robustness of vision models [18].
 Each dataset contains 15 distinct corruption types as tasks (e.g., Gaussian noise, Shot noise, …).
@@
 The text features $\mathcal{Z}$, in zero-shot evaluation, are always pre-computed.
-We also draw a comparison with CLIP performance on respective source test sets and follow  [^1] https://github.com/LAION-AI/CLIP_benchmark for this implementation.
-Throughout this paper, the same TTA problem setup is employed.
+We also draw a comparison with CLIP performance on respective source test sets and follow this implementation  [^1] https://github.com/LAION-AI/CLIP_benchmark.
+Throughout this paper, the same TTA problem setup is used, unless mentioned otherwise.
+This TTA setup differs from the TTA setups of [36, 15, 34].
 Figure 2: Task-wise mean accuracy (%) of zero-shot CLIP across corruption severity levels.
@@
 [Bottom]: ViT-B/16 backbone.
-Dashed lines show zero-shot CLIP performance on corresponding source test sets.
-Prompt Template Backbone
-CIFAR-10C
-CIFAR-100C
-ImageNet-C
-“a photo of a ¡CLS¿.”
-RN101
-42.30 (-36.50) 20.62 (-25.48) 14.97 (-46.23)
-ViT-B/16
-61.16 (-28.84) 35.78 (-30.82) 24.51 (-42.16)
-ViT-B/32
-59.00 (-29.20) 31.79 (-30.51) 23.20 (-38.80)
-ViT-L/14
-75.84 (-19.36) 47.82 (-27.78) 39.55 (-33.95)
-“a bad photo of a ¡CLS¿.”
-RN101
-42.73 (-36.47) 20.18 (-27.02) 12.87 (-49.43)
-ViT-B/16
-62.97 (-28.03) 36.37 (-30.33) 25.84 (-42.26)
-ViT-B/32
-60.25 (-29.75) 31.30 (-31.00) 23.58 (-39.22)
-ViT-L/14
-76.08 (-19.22) 48.09 (-25.41) 39.37 (-34.73)
-“a blurry photo of a ¡CLS¿.”
-RN101
-45.32 (-34.98) 20.28 (-26.12) 15.80 (-45.70)
-ViT-B/16
-62.49 (-28.31) 35.17 (-30.83) 25.27 (-42.03)
-ViT-B/32
-57.85 (-30.65) 31.47 (-30.85) 23.40 (-38.80)
-ViT-L/14
-73.77 (-20.73) 48.07 (-25.73) 39.22 (-34.18)
-“a noisy photo of a ¡CLS¿.”
-RN101
-44.56 (-34.34) 20.71 (-25.89) 15.48 (-45.72)
-ViT-B/16
-63.03 (-27.87) 35.14 (-30.76) 25.05 (-41.35)
-ViT-B/32
-59.76 (-29.24) 31.53 (-30.37) 23.11 (-38.19)
-ViT-L/14
-75.36 (-19.54) 47.87 (-27.63) 38.78 (-34.42)
-Table 1: Mean classification accuracy (%) across all corruption tasks with various prompt templates. We report the accuracy drop (-) (%) relative to performance on the corresponding source test set.
+Dashed lines show zero-shot CLIP performance on corresponding source test sets (clean).
 3.1 Sensitivity of CLIP to image corruption severity
 We analyze CLIP’s zero-shot performance by progressively increasing corruption severity for RN101 and ViT-B/16 vision backbones.
-We report mean accuracy and overall performance in Figure 2.
+We illustrate mean accuracy and overall performance in Figure 2.
 Despite CLIP’s robust multimodal feature space, accuracy drops significantly with an increase in corruption severity, regardless of the backbone.
@@
 This results in lower likelihoods and thus, higher misclassification rates.
+(a) CIFAR-10C
+(b) CIFAR-100C
+(c) ImageNet-C
+Figure 3: Mean classification accuracy (in %) across all the corruption types vs. accuracy on corresponding source test sets, evaluated using various prompt templates and across vision backbones.
 3.2 Sensitivity of CLIP to prompt templates
@@
 Each prompt adds context to help CLIP extract more relevant text features.
-We report the mean classification accuracy (in %) across RN101, ViT-B/16, ViT-B/32, and ViT-L/14 backbones at an image corruption severity level of 5 for all datasets, with results summarized in Table 1.
-The absolute accuracy drop compared to zero-shot performance on the source test set is also reported.
+We report the mean classification accuracy (in %) across RN101, ViT-B/16, ViT-B/32, and ViT-L/14 backbones at an image corruption severity level of 5 for all datasets, with results presented in Figure 3.
+For each dataset, the x-axis is the source accuracy and y-axis is the mean accuracy, across 15 tasks.
 Analysis.
@@
 As expected, RN101 performs worse than the ViT-based backbones, primarily due to its lack of global attention-based modeling inherent to transformers [50].
-Therefore, for the remainder of the experiments in this paper, we focus on ViT-based backbones, specifically ViT-B/16 and ViT-B/32.
-Due to GPU memory limitations, we do not use ViT-L/14 for model adaptation.
+Therefore, for the remainder of the experiments in this paper, we focus on ViT-based backbones, specifically ViT-B/16.
 Unsuitability of prompt template selection at test-time.
 Additionally, at test-time, it is impractical to perform prompt engineering or optimize prompt vectors since 1) Choosing different prompt templates for generating text features is extremely tedious and time-consuming.
+As discussed in Section 2, CLIP TTA works [36, 15, 34] use various prompt templates, making it unrealistic.
 2) In real-time deployment involving prompt-tuning, such prompts cannot quickly estimate the distribution of incoming test batches.
 TPT [45] optimizes pre-trained text prompts for each test image, turning out to be suboptimal since the prompts are optimized ignorant of the distribution of the test image.
-4 BAT-CLIP
-Figure 3: BAT-CLIP not only adapts the visual encoder for highly discriminative image features but also promotes a strong alignment between image and text features by adapting the text encoder too, leading to improved performance following test-time adaptation.
+4 Proposed Methodology
+Figure 4: Illustration of our proposed approach: BATCLIP not only adapts the visual encoder for highly discriminative image features but also promotes a strong alignment between image and text features by adapting the text encoder too, leading to improved performance following test-time adaptation.
 We adapt only the LayerNorm parameters of CLIP encoders.
@@
 Similarly, the performance of the text encoder $f_{txt}$ is invariant to the different text prompt templates and is also impractical to tune at test-time.
-Therefore, for effective adaptation, both the image and text features of CLIP need to be adapted to the incoming domain of test batches.
-We illustrate our proposed framework, BAT-CLIP, in Figure 3.
+Therefore, for effective adaptation, both the image and text encoders of CLIP need to be adapted to the incoming domain of test batches.
+We illustrate our proposed framework, BATCLIP, in Figure 4.
 From unimodal adaptation to bimodal adaptation.
@@
 Projection matching between the visual and text features.
-We propose learning visual and text features that are domain-specific and mutually aware by jointly updating the encoders, as in Figure 3 (middle row), opposed to a unimodal update in TPT.
+We propose learning visual and text features that are domain-specific and mutually aware by jointly updating the encoders, as in Figure 4 (middle row), opposed to a unimodal.
 From Eq.
 1, the visual feature $v_{k,i}^{t}$ of the $k^{th}$ image in batch $x_{i}^{t}$ needs to have a high similarity with the text feature $z_{c}$ of class c for good alignment.
-To quantify this similarity, a possible direction is to project the visual feature onto the text feature i.e., compute the scalar projection — $v_{k,i}^{t}\cdot\hat{z}_{c}$, where $\hat{z}_{c}$ is the normalized text feature of $\hat{z}_{c}$ i.e., $\hat{z}_{c}=\frac{z_{c}}{||z_{c}||_{2}}$.
+To quantify this similarity, an approach is to project the visual feature onto the text feature i.e., compute the scalar projection — $v_{k,i}^{t}\cdot\hat{z}_{c}$, where $\hat{z}_{c}$ is the normalized text feature of $\hat{z}_{c}$ i.e., $\hat{z}_{c}=\frac{z_{c}}{||z_{c}||_{2}}$.
 Geometrically speaking, maximizing this projection leads to more similarity between the corresponding features and hence, a better alignment.
-A possible approach would be to compute the projection for each class-specific visual feature in a test batch.
-However, within the batch, due to the image corruption, the pseudo-labels could be wrong/noisy.
+However, computing such a projection could have limitations.
+Within the test batch, due to the image corruption, the pseudo-labels could be wrong/noisy.
 Instead, we propose modeling the projection of the class prototype with its corresponding text feature.
@@
 […]
-i.e., we compute the mean feature of all the support visual features constituting a class c, based on the computed pseudo-labels (Eqs.
+i.e., for the current test batch, we compute the mean feature of all the support visual features constituting a class c, based on the computed pseudo-labels (Eqs.
 1 and 2), where $\hat{y}$ refers to the pseudo-labels computed via Eq.
@@
 3, class separation, in addition, is needed for good generalization, robustness, and adaptation.
-So, we increase the cosine distance between the class prototypes and enhance the discriminative nature as,
+Therefore, we additionally incorporate a loss to increase the cosine distance between the class prototypes,
 $\mathcal{L}_{sp}=\sum_{l\in C}\sum_{c\in C}\mathds{1}[l\neq c](1-\text{cos}(\bar{v}_{c},\bar{v}_{l}))$ (4)
@@
 $\underset{\phi_{v},\phi_{t}}{\arg\!\min}(\mathcal{L}_{ent}-\mathcal{L}_{pm}-\mathcal{L}_{sp})$ (6)
-where $\phi_{v}$ and $\phi_{t}$ refer to the parameters of $f_{vis}$ and $f_{txt}$, respectively.
+where $\phi_{v}$ and $\phi_{t}$ refer to the LayerNorm parameters of $f_{vis}$ and $f_{txt}$, respectively.
 Motivation of choice of parameters.
-Through thorough experimentation, [50] demonstrates that fine-tuning LayerNorm parameters in attention-based models leads to yielding strong results on multimodal tasks.
+Through thorough experimentation, Zhao et al.
+[58] demonstrate that fine-tuning LayerNorm (LN) parameters in attention-based models leads to yielding strong results on multimodal tasks.
 This also brings a reduction in trainable parameters.
@@
 For every new task/corruption, we reset the model parameters following TENT [53] i.e., we perform a single domain TTA.
-We use a generic hand-crafted prompt template “a photo of a ¡CLS¿.” at all times, thus avoiding the need for a prompt search.
-In summary, our bimodal test-time adaptation approach jointly updates the LayerNorm parameters of the CLIP encoders that are optimized synergistically through loss components aware of the input domain, leading to a more robust multimodal learning process.
+In summary, our bimodal online test-time adaptation approach jointly updates the LayerNorm parameters of the CLIP encoders that are optimized synergistically through loss components aware of the input domain, leading to a more robust multimodal learning process.
 5 Experiments and Results
-5.1 Online Test-Time Adaptation
 Baselines.
-We compare our approach with zero-shot CLIP and other proposed online TTA methods using CLIP - TPT [38] and VTE [9]  [^2] We do not benchmark against WATT [29] since they update the model several times, for each prompt template, over a number of prompt templates..
-For TPT, 63 random augmentations are generated for a single test image in the vision space.
-Based on a threshold, the marginal entropy of confident predictions is minimized to optimize the prompts.
-In the text space, VTE considers an ensemble of hand-crafted prompts for a test image, to get the final predictions (Eq.
-1).
-Alongside this, they also consider multiple random augmentations of the test image, with frozen CLIP encoders.
-We also contrast our work against state-of-the-art TTA methods i.e., TENT [46], RoTTA [47], RPL [34], and SAR [28].
-In essence, we adopt these approaches to use CLIP.
-Following the guidelines set by [9], we update only the vision encoder $f_{vis}$ for such a setup.
+We compare our approach with zero-shot CLIP (ZS), state-of-the-art online TTA methods adopted for CLIP following [10] i.e., TENT [53], RoTTA [54], RPL [41], and SAR [35].
+We update only the vision encoder $f_{vis}$ for such a setup.
+We benchmark against other online TTA methods using CLIP - TPT [45] and VTE [10].
+Based on the discussion in Section 2, we adapt the two variants of WATT [36], for our setup, to WATT-P* and WATT-S*, while maintaining method consistency.
+For a “fairer” comparison, on each test batch, the model is adapted once using each of the suggested 8 prompt templates.
+Note that this setup is not online yet.
+In addition, CLIP parameters are reset after every task.
+We report the results on CIFAR-10C and CIFAR-100C only, based on their suggested settings.
+On ImageNet-C, we also report results against online StatA [55], based on their chosen hyperparameters.
 The model predictions can then be computed as in Eq.
 1.
-We mention the training details of each TTA method in the Supplementary.
+We mention the details of each TTA method in the Supplementary.
 Implementation Details.
-We query the text encoder $f_{txt}$ with a general and fixed hand-crafted prompt template “a photo of a ¡CLS¿.” for all the datasets, as motivated earlier.
+We query the text encoder $f_{txt}$ with a general prompt template “a photo of a $<$CLS$>$.” for all the datasets, as motivated earlier.
 We optimize both the vision encoder $f_{vis}$ and text encoder $f_{txt}$.
@@
 Similarly, for CIFAR-100C and ImageNet-C, Adam and AdamW optimizers are respectively used, at a fixed learning rate of 5$\times$$10^{-4}$, with the model being reset after each task.
-The batch sizes are set to 200, 200, and 64 for the datasets, following various TTA benchmarks, at a corruption severity level of 5 for each task.
-As visual backbones, we report results on ViT-B/16 and ViT-B/32 [10], where all of our experiments are run on a single NVIDIA RTX A5000 GPU.
-Method
+For fairness to all the baselines, the batch sizes are set to 200, 200, and 64 for the datasets, following various TTA benchmarks, at a corruption severity level of 5 for each task.
+For WATT-P* and WATT-S*, we use the same templates and learning rates as mentioned in [36].
+For StatA, the class correlation in a batch is controlled by a Dirichlet distribution ($\gamma$).
+We report results for $\gamma$ = 0.1 (low correlation) and -1 (sequential).
+All experiments are run on a single NVIDIA RTX A5000 GPU using ViT-B/16 as the visual backbone.
+Results on ViT-B/32 are reported in the Supplementary.
+Method Venue
 Gaussian
@@
 CIFAR-10C
-ViT-B/16
-37.92 41.7 54.42 71.75 40.89 67.93 73.62 73.89 77.35 70.22 84.45 62.36 53.81 47.65 59.43 61.16
+ZS
+ICLR’21 37.92 41.7 54.42 71.75 40.89 67.93 73.62 73.89 77.35 70.22 84.45 62.36 53.81 47.65 59.43 61.16
+TENT
+ICLR’21 15.49 18.28 38.12 81.59 21.73 76.32 82.35 84.62 82.19 80.60 91.83 80.55 63.52 58.57 54.71 62.03
+RoTTA
+CVPR’23 39.17 42.90 55.41 72.18 41.32 68.02 74.01 74.38 78.01 70.80 84.80 63.19 54.62 49.32 60.15 61.89
+RPL
+arXiv 15.47 17.43 40.73 81.76 20.08 69.89 82.93 84.43 83.19 81.84 91.80 79.42 64.89 54.07 54.90 61.52
+SAR
+ICML’22 47.98 53.60 60.56 74.30 47.56 73.15 76.43 77.91 79.88 75.66 86.79 71.62 58.34 62.03 64.71 67.37
+[…]
 TPT
-37.74 42.24 60.57 72.88 44.80 69.69 75.37 75.96 78.84 72.12 85.68 62.04 58.90 55.14 62.64 63.64
+NeurIPS’22 37.74 42.24 60.57 72.88 44.80 69.69 75.37 75.96 78.84 72.12 85.68 62.04 58.90 55.14 62.64 63.64
 VTE
-42.42 46.26 64.23 71.10 45.58 68.50 73.66 76.75 78.27 71.02 85.28 57.24 59.54 60.59 61.85 64.15
+ECCV-W’24 42.42 46.26 64.23 71.10 45.58 68.50 73.66 76.75 78.27 71.02 85.28 57.24 59.54 60.59 61.85 64.15
+WATT-P*
+NeurIPS’24 44.70 49.66 60.48 74.35 46.33 72.21 76.28 78.18 80.97 74.90 87.66 68.63 58.35 56.13 63.47 66.15
+WATT-S*
+NeurIPS’24 57.22 61.86 67.63 78.61 53.07 77.85 80.14 81.84 83.46 80.38 89.67 78.45 65.76 68.63 67.67 72.81
 Ours
 61.13 64.09 65.76 80.51 54.96 80.65 81.94 83.04 84.19 80.84 88.95 82.15 69.16 62.68 66.64 73.85
-Gain/Loss(%)
-+18.71 +17.83 +1.53 +7.63 +9.38 +10.96 +6.57 +6.29 +5.35 +8.72 +3.27 +19.79 +9.62 +2.09 +4.00 +9.70
-ViT-B/32
-35.47 39.94 43.23 69.95 41.43 64.50 70.13 70.85 72.33 66.66 81.37 64.57 59.69 48.28 56.62 59.00
+CIFAR-100C
+ZS
+ICLR’21 19.64 21.40 25.26 42.54 20.03 43.17 47.95 48.35 49.74 41.57 57.02 34.58 29.15 23.96 32.43 35.79
+TENT
+ICLR’21 7.60 8.21 8.33 51.81 7.95 52.45 55.34 54.16 36.17 50.92 65.63 54.51 36.52 43.99 35.81 37.96
+RoTTA
+CVPR’23 20.65 22.22 26.17 42.48 20.26 42.90 47.88 48.75 49.92 41.86 57.00 34.52 29.27 25.08 32.88 36.12
+RPL
+arXiv 6.44 7.09 7.09 52.16 11.81 52.33 55.50 54.20 38.83 51.99 66.07 54.45 36.86 42.83 39.45 38.47
+SAR
+ICML’22 25.30 27.19 32.78 47.12 23.42 47.16 51.70 51.94 52.48 48.77 61.54 44.50 32.26 33.67 38.06 41.19
+[…]
 TPT
-43.11 46.53 48.29 71.31 47.80 66.89 71.96 74.00 76.00 68.81 84.12 66.35 63.86 51.86 58.01 62.59
+NeurIPS’22 17.95 19.51 27.13 43.53 20.08 42.65 48.63 49.11 49.48 42.14 57.35 33.26 31.13 27.59 32.75 36.15
 VTE
-47.59 50.18 53.15 71.39 53.86 67.92 72.90 76.37 76.30 70.78 83.27 61.07 69.00 58.57 61.14 64.90
-Ours
-52.39 55.99 52.54 76.79 54.04 74.90 75.79 77.67 79.10 75.31 86.33 77.34 67.41 57.06 61.29 68.26
-Gain/Loss(%)
-+4.80 +5.81 -0.61 +5.40 +0.18 +6.98 +2.89 +1.30 +2.80 +4.53 +2.21 +10.99 -1.59 -1.51 +0.15 +3.36
-CIFAR-100C
-ViT-B/16
-19.64 21.40 25.26 42.54 20.03 43.17 47.95 48.35 49.74 41.57 57.02 34.58 29.15 23.96 32.43 35.79
-TPT
-17.95 19.51 27.13 43.53 20.08 42.65 48.63 49.11 49.48 42.14 57.35 33.26 31.13 27.59 32.75 36.15
-VTE
-17.96 18.72 28.17 40.38 19.60 39.50 45.33 48.24 46.87 40.73 55.31 30.04 32.47 30.35 31.45 35.01
+ECCV-W’24 17.96 18.72 28.17 40.38 19.60 39.50 45.33 48.24 46.87 40.73 55.31 30.04 32.47 30.35 31.45 35.01
+WATT-P*
+NeurIPS’24 20.53 22.22 27.3 43.14 17.51 42.37 48.17 47.31 49.34 41.49 57.07 35.29 27.75 25.83 31.89 35.81
+WATT-S*
+NeurIPS’24 21.20 23.11 28.23 44.16 18.45 43.44 49.14 48.16 50.05 42.35 57.82 36.43 28.36 26.85 32.93 36.71
 Ours
 24.91 27.73 33.66 50.11 26.27 48.49 54.85 52.35 51.62 48.38 63.27 45.21 34.74 32.38 37.31 42.09
-Gain/Loss(%)
-+5.27 +6.33 +5.49 +6.58 +6.19 +5.32 +6.22 +3.24 +1.88 +6.24 +5.92 +10.63 +2.27 +2.03 +4.56 +5.94
-ViT-B/32
-16.23 17.83 17.57 39.07 17.63 38.55 43.81 42.32 43.46 39.71 50.32 29.34 28.74 22.85 29.42 31.79
+ImageNet-C
+ZS
+ICLR’21 11.18 12.54 12.04 23.36 15.18 24.50 22.58 32.32 29.88 35.88 54.18 17.20 12.72 30.96 33.26 24.51
+TENT
+ICLR’21 5.14 5.70 7.44 25.22 19.34 26.80 24.16 33.56 30.42 37.74 54.24 22.50 13.90 35.02 36.08 25.15
+RoTTA
+CVPR’23 11.34 12.96 12.32 23.38 15.50 24.66 22.90 32.56 30.02 35.98 54.32 17.20 12.80 31.06 33.46 24.78
+RPL
+arXiv 9.04 10.04 10.96 24.40 17.40 26.28 23.76 32.70 30.62 36.64 54.04 19.38 13.24 33.14 34.60 25.08
+SAR
+ICML’22 17.96 20.46 20.68 25.72 23.04 29.52 26.04 34.92 32.74 39.00 55.00 27.14 19.64 36.66 37.50 29.73
+[…]
 TPT
-16.08 17.65 17.54 39.21 19.47 38.91 44.01 43.45 44.46 40.15 50.93 27.77 30.91 23.36 29.55 32.23
+NeurIPS’22 8.48 9.46 10.20 23.98 15.16 25.10 24.00 33.94 32.12 37.08 55.64 16.54 13.68 34.06 33.58 24.87
 VTE
-16.84 18.33 18.94 39.63 22.88 39.13 43.80 44.56 44.88 39.21 49.37 28.37 34.13 26.87 30.12 33.14
-Ours
-21.35 24.71 22.32 46.26 23.07 44.64 50.12 47.23 46.88 44.92 58.55 38.52 34.56 27.73 33.19 37.60
-Gain/Loss(%)
-+4.51 +6.38 +3.38 +6.63 +0.19 +5.51 +6.11 +2.67 +2.00 +4.77 +7.62 +9.18 +0.43 +0.86 +3.07 +4.46
-ImageNet-C
-ViT-B/16
-11.18 12.54 12.04 23.36 15.18 24.50 22.58 32.32 29.88 35.88 54.18 17.20 12.72 30.96 33.26 24.51
-TPT
-8.48 9.46 10.20 23.98 15.16 25.10 24.00 33.94 32.12 37.08 55.64 16.54 13.68 34.06 33.58 24.87
-VTE
-9.18 10.76 10.78 24.72 14.30 24.36 25.24 35.38 32.46 38.16 55.56 16.14 14.26 38.72 33.98 25.60
+ECCV-W’24 9.18 10.76 10.78 24.72 14.30 24.36 25.24 35.38 32.46 38.16 55.56 16.14 14.26 38.72 33.98 25.60
+StatA ($\gamma$=0.1)
+CVPR’25 10.56 11.22 10.86 22.11 14.12 22.39 20.85 31.39 29.90 34.63 53.32 16.00 12.46 29.55 32.79 23.47
+StatA ($\gamma$=-1)
+CVPR’25 11.83 13.02 12.41 23.71 15.02 23.64 21.79 31.80 30.22 36.52 54.15 17.56 13.00 32.01 33.37 24.67
 Ours
 19.32 21.38 19.60 26.58 21.94 30.88 29.02 36.48 32.00 40.98 56.72 26.14 23.74 37.67 38.34 30.72
-Gain/Loss(%)
-+8.14 +8.84 +7.56 +1.86 +6.76 +5.78 +3.78 +1.10 -0.46 +2.82 +1.08 +8.94 +9.48 -1.05 +4.36 +5.12
-ViT-B/32
-12.88 13.04 12.90 24.42 11.86 22.72 20.20 25.70 25.84 30.28 50.54 17.32 18.96 32.20 29.12 23.20
-TPT
-12.04 12.64 12.52 25.38 12.28 22.68 20.78 26.36 26.64 30.78 51.02 16.50 19.90 33.62 30.62 23.58
-VTE
-11.96 12.32 13.44 25.06 11.70 22.58 22.40 27.38 27.02 32.28 51.52 16.84 19.94 34.80 32.82 24.14
-Ours
-16.84 18.20 16.10 25.04 20.90 28.90 25.24 29.42 27.18 36.02 50.18 17.66 27.68 36.20 35.42 27.39
-Gain/Loss(%)
-+3.96 +5.16 +2.66 -0.34 +8.62 +6.18 +2.84 +2.04 +0.16 +3.74 -1.34 +0.34 +7.74 +1.40 +2.60 +3.25
-Table 2: Mean accuracy (%) on CIFAR-10C, CIFAR-100C, and ImageNet-C - TTA mean accuracy of the 15 corruptions (tasks) at a severity level of 5, using ViT-B/16 and ViT-B/32.
-+ and - denote the absolute gain/loss w.r.t the next best performance.
-We contrast our results against zero-shot ViT-B/16 and -B/32, TPT [38], and VTE [9].
+Table 1: Mean accuracy (%) on CIFAR-10C, CIFAR-100C, and ImageNet-C - TTA mean accuracy of the 15 corruptions (tasks) at a severity level of 5, using ViT-B/16.
+5.1 Results: Online Test-Time Adaptation
+Results on CIFAR-10C, CIFAR-100C, and ImageNet-C.
+We present the online TTA results in Table 1.
+While most TTA methods outperform zero-shot CLIP on average, BATCLIP consistently exceeds the performance of these methods across all datasets.
+SAR shows significant performance improvements over others.
+Although it performs similarly to our method on CIFAR-100C and ImageNet-C, BATCLIP outperforms SAR on CIFAR-10C, achieving improvements of 6.48%.
+While all the prior approaches operate solely in the vision space, our approach leverages CLIP’s joint vision-language feature space, resulting in superior TTA performance.
+For TTA approaches using CLIP, BATCLIP surpasses TPT and VTE across all tasks and datasets.
+For our WATT-P* and WATT-S* implementation, adaptation happens once on each prompt template for a batch.
+Interestingly, despite CLIP being trained on the batch by averaging adapted weights from multiple templates before inference, the performance remains subpar.
+This indicates that WATT [36] is not well suited for online TTA.
+In StatA [55], $\gamma$ controls the number of effective classes in a batch, for zero-shot VLM adaptation.
+On ImageNet-C, the large drop in performance for each task indicates the need for CLIP adaptation to severe image degradation and the importance of relevant text features.
+In Figure 5, we show t-SNE [49] plots of visual and text features, of a few tasks, and compare them against zero-shot ViT-B/16 for CIFAR-10C.
+It highlights how our approach enhances alignment with text features, fosters better class separation, and forms more distinct clusters, leading to significant improvements.
+The Supplementary shows detailed t-SNE plots for CIFAR-10C and CIFAR-100C.
 Gaussian Defocus Fog Pixelate
@@
 Ours Refer to caption Refer to caption Refer to caption Refer to caption
-Figure 4: BAT-CLIP yields more discriminative visual features that exhibit stronger alignment with their corresponding text features.
+Figure 5: BATCLIP yields more discriminative visual features that exhibit stronger alignment with their corresponding text features.
 The t-SNE [49] plots show visual ($\circ$) and text ($\bigstar$) features for CIFAR-10C, comparing zero-shot ViT-B/16 with our approach.
-Results on CIFAR-10C, CIFAR-100C, and ImageNet-C.
-We present the TTA results in Table 2.
-Our method significantly improves the mean accuracy across all tasks and datasets for both visual backbones.
-Unlike TPT [38] and VTE [9], which process a single test image at a time and therefore incur substantial time, our approach operates at the batch level.
-This allows us to leverage the semantic relationships within the batch.
-By updating the parameters of $f_{vis}$ and $f_{txt}$ through the projection matching loss and enhancing the cosine distance between the class prototypes, our method effectively strengthens the alignment between visual and text features and learns discriminative visual features.
-In Figure 4, we show t-SNE [43] plots of visual and text features, of a few tasks, and compare them against zero-shot ViT-B/16 for CIFAR-10C.
-It highlights how our approach enhances alignment with text features, fosters better class separation, and forms more distinct clusters, leading to significant improvements compared to zero-shot CLIP.
-The Supplementary shows detailed t-SNE plots for CIFAR-10C and CIFAR-100C.
-Concerning compute time per task, on a difficult dataset like ImageNet-C, our method takes about 45 s, compared to 40 mins for TPT and 4 mins for VTE.
-This makes our approach deployment-friendly for quick classification with large improvements.
-Dataset Backbone Zero-Shot TENT RoTTA RPL SAR Ours
-CIFAR-10C ViT-B/16 61.16 62.06 61.89 61.52 67.37 73.85
-ViT-B/32 59.00 56.35 59.57 56.51 65.13 68.26
-CIFAR-100C ViT-B/16 35.79 37.96 36.12 38.47 41.19 42.09
-ViT-B/32 31.79 31.61 31.84 30.00 37.81 37.60
-ImageNet-C ViT-B/16 24.51 25.15 24.78 25.08 29.73 30.72
-ViT-B/32 23.20 24.05 23.33 23.62 28.07 27.39
-Table 3: Mean accuracy (%) of CLIP adopted for TTA approaches at a corruption severity level of 5, using ViT-B/16 and ViT-B/32.
-Results against prior TTA methods.
-In Table 3, we compare our method with existing TTA approaches, all adopted for CLIP.
-We report mean accuracy across all tasks.
-While most TTA methods outperform zero-shot CLIP on average, our approach consistently matches or exceeds the performance of these methods across all datasets and visual backbones.
-Detailed per-task accuracies are provided in the Supplementary.
-With a ViT-B/16 backbone, TENT [46] increases the accuracy over zero-shot CLIP by 0.9%, 2.18%, and 0.64%, for the respective benchmark datasets.
-SAR [28] shows significant performance improvements across both vision backbones over prior TTA methods.
-Although it performs similarly to our method on most benchmarks, our approach outperforms SAR on CIFAR-10C, achieving improvements of 6.48% and 3.13% on the ViT-B/16 and ViT-B/32 backbones, respectively.
-While all the prior approaches operate solely in the vision space, our approach leverages CLIP’s joint vision-language feature space, resulting in superior TTA performance.
+Dataset Domain ZS TENT WATT-P* WATT-S* Ours
+OfficeHome Art 78.29 79.4 80.35 80.43 79.86
+Clipart 64.03 63.18 66.85 66.9 66.44
+Product 87.11 88.42 87.5 87.54 88.51
+Real World 88.96 89.6 90.04 89.99 89.67
+PACS Art 97.22 98.05 97.75 97.75 97.56
+Cartoon 97.4 97.65 97.53 97.53 97.48
+Photo 99.58 99.58 99.59 99.52 99.72
+Sketch 86.23 88.75 88.52 88.65 87.76
+VLCS Caltech101 99.43 99.43 99.36 99.36 99.51
+LabelMe 68.15 68.14 66.92 68.49 68.94
+SUN09 73.4 73.4 74.53 74.68 75.23
+VOC2007 84.75 84.75 84.0 84.03 85.6
+TerraInc L38 20.3 20.3 27.79 29.08 37.33
+L43 31.52 31.52 33.98 34.13 32.71
+L46 28.98 28.98 27.07 28.13 31.05
+L100 52.35 52.35 43.59 42.32 55.22
+Table 2: Accuracy on domain generalization datasets [14] with other domain shifts using a ViT-B/16 visual backbone.
+5.2 Results: Domain Generalization Datasets
+In this section, we evaluate our BATCLIP on commonly used domain generalization datasets [14] and benchmark against zero-shot ViT-B/16 (ZS), TENT, WATT-P* and WATT-S*.
+From Table 2, we observe that BATCLIP does better or comparable, across all datasets and domain shifts, to all the baselines.
+TENT matches zero-shot CLIP since LN normalization works per image across the features, independent of the batch statistics.
+We demonstrate that BATCLIP can be generalized and be robust to other various domain shifts, with a single step adaptation and global prompt template as opposed to WATT-P* and WATT-S* that use multiple prompt templates.
+The largest performance improvements come on the Terra Incognita dataset.
+The implementation details are described in the Supplementary.
 5.3 Ablation Study and Analysis
@@
 Adaptation for multiple iterations.
-We adapt CLIP at test-time with one update per batch.
-Here, we analyze adaptation for multiple iterations on a single batch.
-In Figure 5, we illustrate the mean accuracy, across all 15 tasks, using the ViT-B/16 backbone.
-Continuous adaptation of LayerNorm parameters, which normalize along the features of a single sample, can lead to overfitting as it fits to the noise patterns of the batch, ultimately degrading generalization to other domains.
+Here, we analyze adaptation for multiple iterations on a single batch, shown in Figure 6, using the ViT-B/16 backbone.
+Continuous adaptation of LN parameters, which normalize along the features of a single sample, can lead to overfitting as it fits to the noise patterns of the batch, ultimately degrading generalization to other domains.
 This could also lead to a decline in the loss of CLIP pre-trained knowledge.
-In the Supplementary, we report the post-adaptation results back on the source test sets - CIFAR10 and CIFAR100.
-Our approach still maintains a significant advantage over zero-shot CLIP evaluation.
-Despite a gradual decline in mean accuracy, the performance gap remains notably better, particularly on CIFAR-10C and ImageNet-C.
-This suggests that the proposed loss components enhance robustness, maintaining effectiveness under challenging conditions.
+In the Supplementary, we report the post-adaptation results back on the source test sets - CIFAR10 and CIFAR100 to evaluate catastrophic forgetting.
+Our approach outperforms zero-shot CLIP, with a smaller decline in mean accuracy—especially on CIFAR-10C and ImageNet-C—showing that our loss components improve robustness under challenging conditions.
 Backbone (ViT-B/16) CIFAR-10C CIFAR-100C ImageNet-C
@@
 Table 3: Ablation on loss components - Mean accuracy (in %).
-Contribution of each loss component.
-In Table 4, we validate the effectiveness of the loss components by taking their different combinations.
-The mean accuracies across all the tasks are reported using ViT-B/16.
+Ablation on loss components.
+In Table 3, we ablate the loss components by taking their different combinations using ViT-B/16.
 The task-wise results are reported in the Supplementary.
@@
 Interestingly, using just $\mathcal{L}_{ent}$ achieves better or comparable accuracy than zero-shot CLIP, as the encoders work in synergy while adapting to the domain.
-Impact of bimodal adaptation.
-To show the effectiveness of bimodal adaptation, we ablate $\mathcal{L}_{pm}$ in Eq.
-6, which is responsible for updating the text encoder $f_{txt}$.
-We report the mean classification results using the ViT-B/16 vision backbone in Table 5.
-As can be seen, we see a drop in accuracy when $f_{txt}$ is “frozen” i.e., when $\mathcal{L}_{pm}$ isn’t used, necessitating the need for bimodal adaptation of CLIP encoders.
-Dataset $f_{vis}$ update $f_{txt}$ update Ours
-CIFAR-10C 72.58
-73.85
-CIFAR-100C 41.00
-42.09
-ImageNet-C 29.88
-30.72
-Table 5: Ablation on $\mathcal{L}_{pm}$ to demonstrate the need of bimodal adaptation of CLIP encoders - using a ViT-B/16 backbone.
 Refer to caption Refer to caption
-Figure 6: Ablation on low batch size - CIFAR-10C and CIFAR-100C (left) and ImageNet-C (right) using ViT-B/16 backbone.
+Figure 7: Ablation on low batch size - CIFAR-10C and CIFAR-100C (left) and ImageNet-C (right) using a ViT-B/16 backbone.
 Low batch size setting.
-In Eq 3, the prototype of each class, via pseudo-labels of each test batch, can be affected by the batch size, with a lower size indicating the absence of a few classes.
+In Eq 3, the prototype of each class can be affected by the batch size, with a lower size indicating the absence of a few classes.
 In Figure 7, we illustrate the mean accuracies on the benchmark datasets.
-As seen, BAT-CLIP still outperforms or performs comparably to TPT and VTE in low batch size settings.
-Since TTA happens at the batch-level of a specific domain, the prototypes are computed with the available pseudo-labels, aiding in the continual learning of the feature space as new batches come by.
-Effect of different prompt templates.
-In Table 6, we show results with “relevant” prompt templates to show the independence of such prompt selection, at test-time.
-As seen, the performance gain over zero-shot ViT-B/32 is fairly large for all the prompt templates.
-Though TPT fine-tunes a pre-trained prompt on each image, and VTE uses an ensemble of prompts, our method is agnostic to the prompt template being used, making it favorable for real-time usage.
-Prompt Template CIFAR-10C CIFAR-100C ImageNet-C
-“a low contrast photo of a ¡CLS¿.”
-68.53 (+7.81) 37.09 (+4.97) 27.31 (+3.71)
-“a blurry photo of a ¡CLS¿.”
-68.84 (+10.96) 36.80 (+5.33) 26.92 (+3.52)
-“a photo of a big ¡CLS¿.”
-67.49 (+10.10) 35.79 (+4.87) 25.64 (+3.29)
-Table 6: Prompt template selection. + denotes the accuracy gain over zero-shot ViT-B/32.
+As seen, BATCLIP still outperforms or performs comparably to TPT and VTE in low batch size settings.
+Since TTA happens at the batch-level of a specific domain, the prototypes are computed with the available pseudo-labels, aiding in the continual learning of the feature space as new batches come.
+Dataset ZS TPT VTE Ours
+CIFAR-10C 75.84 77.75 75.32 84.74
+CIFAR-100C 47.82 49.62 48.53 49.25
+ImageNet-C 39.55 40.96 40.55 42.84
+(a) Mean acc. w/. a ViT-L/14.
+Dataset TENT RPL RoTTA SAR Ours
+ImageNet-C 40.62 40.00 39.76 42.10 42.84
+(b) TTA baselines vs ours w/ ViT-L/14.
+Table 4: Ablation on a larger backbone (ViT-L/14) - Mean accuracy (in %)
+Extension to larger backbones.
+We run BATCLIP for a larger backbone like ViT-L/14 and report the results in Table 4.
+The key advantage lies in the projection matching loss ($\mathcal{L}_{pm}$) which updates the text encoder to produce more image-aware text features—an aspect missing in other state-of-the-art TTA baselines.
+Computational efficiency.
+Due to an additional text encoder update, BATCLIP updates $\sim 0.044$% of all CLIP parameters, with a ViT-B/16 backbone, as opposed to $\sim 0.026$% in other TTA baselines.
+Though the increase is marginal, this leads to superior performances as adaptation of the text encoder produces input-aware CLIP text features, aligning better with the image features.
+On an NVIDIA RTX A5000 GPU, it takes 0.2s/batch on ImageNet-C with a batch size of 64, compared to 2.34s for WATT [36].
+Also, WATT adapts per template (for several iterations), making it expensive every time.
+BATCLIP achieves faster on-the-fly adaptation rates and is favorable for real-time deployment.
 6 Conclusion
-In this paper, we propose BAT-CLIP, a bimodal online test-time adaptation framework for CLIP aimed at handling diverse image corruptions simulating real-time environments [15].
+In this work, we propose BATCLIP, a bimodal online test-time adaptation framework for CLIP aimed at handling diverse image corruptions simulating real-time environments.
 Our in-depth analysis of CLIP’s zero-shot performance under increasing corruption severity reveals significant shortcomings in generalization, highlighting the need for effective adaptation.
-While prior works on TTA for CLIP have predominantly been unimodal focusing on prompt-tuning [38] or prompt ensemble [9] with no model updates, our approach encourages learning rich class-separated visual features via vision encoder updates and strengthens alignment between the image class prototype and corresponding text feature via the text encoder updates.
-Empirical studies, including ablation results, demonstrate the robustness and substantial performance improvements of our approach over existing methods on benchmark corruption datasets — CIFAR-10C, CIFAR-100C, and ImageNet-C, through synergistic CLIP encoder updates.
+Although prior work on TTA for CLIP has predominantly been unimodal focusing on prompt-tuning or prompt ensemble with no model updates, BATCLIP, through synergistic CLIP encoder updates, encourages learning rich class-separated visual features via vision encoder updates and strengthens alignment between the image class prototype and corresponding text feature via text encoder updates.
+Empirical studies, including ablations, show robust performance gains over all baselines on benchmark corruption and domain generalization datasets.
 Supplementary Material
-In this work, we study the problem of test-time adaptation (TTA) of CLIP [31] towards common image corruptions and propose improved schemes for increasing the robustness of CLIP.
-We put forward a bimodal domain adaptation scheme, wherein we exploit the shared feature space of CLIP.
+In this work, we study the problem of online test-time adaptation (TTA) of CLIP [38] towards common image corruptions and propose improved schemes for increasing the robustness of CLIP.
+In addition, to demonstrate the broad impact of our proposed approach, we evaluate on common domain generalization datasets [14]-OfficeHome [51], PACS [25], VLCS [12], and Terra Incognita [3].
+We put forward a bimodal domain adaptation scheme, at test-time, wherein we exploit the shared feature space of CLIP.
 In essence, leaning towards a more effective multimodal learning and adaptation method, we propose loss components that improve alignment between the class-specific visual prototype and corresponding text features via maximizing the projection.
@@
   1. 1.
-    Subsection 6.1 offers a detailed discussion of the datasets used, supplemented with visual illustrations.
+    Section 6.1 offers a detailed discussion of the standard common corruption datasets [18] used, supplemented with visual illustrations.
   2. 2.
-    To ensure full transparency, we outline the implementation details of all methods in section 6.2, including those for prior TTA approaches [46, 47, 34, 28] adapted for use with CLIP.
+    To ensure full transparency, we outline the implementation details of all the methods in Section 6.2, including those for prior TTA approaches [53, 54, 41, 35] adapted for CLIP.
+We also discuss details of the adapted version of WATT [36] - WATT-P* and WATT-S* for online TTA, and other details of experiments run on the domain generalization datasets.
   3. 3.
@@
     • In subsection 6.3.1, we explore the limitations of zero-shot CLIP when using ViT-B/32 and ViT-L/14 backbones under increasing image corruption severity.
-    • Subsection 6.3.2 expands on task-specific accuracies, by adopting TTA approaches, for CIFAR-10C, CIFAR-100C, and ImageNet-C, along with a detailed performance discussion.
-    • In subsections 6.3.3 and 6.3.4, we present a detailed loss ablation study and the post-adaptation zero-shot generalization on source test sets, respectively — essentially evaluating catastrophic forgetting of BAT-CLIP.
-    • Lastly, we include task-wise t-SNE visualizations in subsection 6.3.5 for CIFAR-10C and CIFAR-100C, comparing our method against zero-shot CLIP (ViT-B/16), to illustrate the effectiveness of BAT-CLIP.
+    • In Subsection 6.3.2, we report the main online TTA results on CIFAR-10C, CIFAR-100C, and ImageNet-C with a ViT-B/32 backbone.
+    • We study the effect of different prompt templates on BATCLIP in Subsection 6.3.4.
+Subsection 6.3.5 discusses the ablation of $\mathcal{L}_{pm}$ which is responsible for updating the text encoder.
+    • In subsections 6.3.3 and 6.3.6, we present a detailed loss ablation study and the post-adaptation zero-shot generalization on source test sets, respectively — essentially evaluating catastrophic forgetting of BATCLIP.
+    • Lastly, we include task-wise t-SNE visualizations in subsection 6.3.7 for CIFAR-10C and CIFAR-100C, comparing our method against zero-shot CLIP (ViT-B/16), to illustrate the effectiveness of BATCLIP.
 6.1 Datasets
@@
 Our approach demonstrates enhanced robustness and higher accuracy, especially in challenging image corruption conditions.
-Method
+6.2 Implementation Details
+In this section, we summarize the implementation details of all the baseline methods that have been mentioned in the main paper, including ours.
+We build our approach on the standard benchmark code base  [^2] https://github.com/mariodoebler/test-time-adaptation/tree/main that also houses the hyperparameters and training details of all the prior TTA methods.
+CLIP-like models are used as provided by OpenCLIP.
+Only the vision encoder is updated for existing online TTA methods [53, 54, 41, 35] adopted for CLIP.
+6.2.1 Experiments on CIFAR-10C, CIFAR-100C, and ImageNet-C
+BATCLIP (Ours): For domain-specific test adaptation, we conducted experiments using ViT-B/16 and ViT-B/32 [11] as the vision backbones.
+For CIFAR-10C, both the vision encoder ($f_{vis}$) and text encoder ($f_{txt}$) were updated using the AdamW optimizer with a learning rate of 10-3.
+Similarly, for CIFAR-100C and ImageNet-C, we employed the Adam optimizer and AdamW optimizer, respectively, with a learning rate of 5$\times$10-4.
+The batch size $\mathcal{B}$ used was set to 200 for CIFAR-10C and CIFAR-100C, and 64 for ImageNet-C.
+Throughout, the prompt template is fixed to “a photo of a $<$CLS$>$.”.
+TENT [53]: We follow all the hyperparameters that TENT provides in their official implementation  [^3] https://github.com/DequanWang/tent.
+To update the vision encoder, we use Adam as the optimizer with a learning rate of 10-3 for CIFAR-10C and CIFAR-100C.
+For ImageNet-C, we update using SGD with a learning rate of 25$\times$10-5.
+RoTTA [54]: For fairness, the batch sizes are set to 200 for the CIFAR datasets and 64 for ImageNet-C.
+The vision encoder is updated based on the Adam rule with a learning rate of 10-3.
+The capacity of the memory bank is set to 64, for all the datasets.
+Following the notations in the paper, $\alpha$ = 0.05, $\delta$ = 0.1, $\nu$ = 0.001, $\lambda_{t}$ and $\lambda_{u}$ = 1.0.
+We implement the details exactly as described in their main paper.
+RPL [41]: We use an Adam optimizer with a learning rate of 10-3 for CIFAR-10C and CIFAR-100C.
+For ImageNet-C, the update rule is SGD with a learning rate of 5$\times$10-4.
+To compute the generalized cross-entropy loss, q is set to 0.8 for all the datasets.
+SAR [35]: The training details/hyperparameters for SAR are the same as RPL [41] for CIFAR-10 and CIFAR-100.
+For ImageNet-C, the learning rate is set to 25$\times$10-5 with an SGD update rule.
+The entropy threshold $E_{0}$ is 0.4xln(C), where C is the number of classes.
+$\rho$ is set to a default of 0.05.
+The moving average factor is 0.9 for $e_{m}$ and $e_{0}$ is set to 0.2.
+All parameters are the same as in [35].
+TPT [45]: For each test image, 63 augmentations are generated based on random resized crops, yielding a batch of 64 images, in addition to the original test image.
+The prompt/context vectors are initialized based on “a photo of a $<$CLS$>$.” and tokenized using pre-trained CLIP weights.
+The confidence threshold is set to 10% i.e., the marginal entropy over the 10% confident samples is minimized.
+For all the datasets, we follow their core implementation and optimize the prompt vectors using an AdamW optimizer with a learning rate of 5$\times$10-3.
+VTE [10]: In VTE, an ensemble of different prompt templates is considered based on the idea of [38].
+An example of templates includes “a photo of a $<$CLS$>$.”, “a sketch of a $<$CLS$>$.”, “a painting of a $<$CLS$>$.”, etc.
+The prompt templates are then averaged.
+On the vision side, similar to TPT [45], a batch of random augmentation is created for a test image with no model updates.
+WATT-P* and WATT-S* [36]: The two original variants of WATT [36] were proposed with weight-averaging of adapted weights from multiple prompt templates.
+Additionally, for each test batch, adaptation was performed over multiple iterations.
+To fit our online TTA scheme, we reduced the number of iterations to a single step for each prompt template and reset CLIP parameters only after a domain.
+However, this is still not fully online, as training is performed on the test batch using 8 selected prompt templates: “a photo of a $<$CLS$>$”, “itap of a $<$CLS$>$”, “a bad photo of the $<$CLS$>$”, “a origami $<$CLS$>$”, “a photo of the large $<$CLS$>$”, “a $<$CLS$>$ in a video game”, “art of the $<$CLS$>$”, and “a photo of the small $<$CLS$>$”.
+As they report performance on CIFAR-10C and CIFAR-100C only, we follow their original implementation details.
+For experiments using WATT-S*, we set a batch size of 200 (for a fair comparison to other baselines) and a learning rate of 10-3 using an Adam optimizer.
+For WATT-P*, a learning rate of 10-4 is used with the same batch size.
+StatA [55]: We adopt the original hyperparameter settings of StatA for online TTA in our reported ImageNet experiments and apply them to ImageNet-C.
+To control the effective number of correlated classes in a test batch, we set $\gamma$ to 0.1 (low correlation) and -1 (separate, sequential).
+The default prompt template is “a photo of a $<$CLS$>$.”
+6.2.2 Experiments on Domain Generalization Datasets
+We evaluate BATCLIP on standard domain generalization datasets [14].
+For a fair comparison, as reported in WATT [36], we use a batch size of 128 for all the online TTA experiments on VLCS, PACS, and Office Home.
+We use an AdamW optimizer for model updates using BATCLIP with learning rates of 5$\times$10-4, 10-3, and 5$\times$10-3 for OfficeHome, PACS, and VLCS, respectively.
+We use the same learning settings for TENT, WATT-S*, and WATT-P*, as in WATT [36].
+In addition to the mentioned datasets, we also run experiments on the Terra Incognita dataset and optimize using an AdamW optimizer with a learning rate of 5$\times$10-4 and a batch size of 256.
+Figure 10: Task-wise mean accuracy (%) of zero-shot CLIP across different corruption severity levels.
+[Top]: ViT-B/32 backbone.
+[Bottom]: ViT-L/14 backbone.
+The dashed lines indicate the performance of zero-shot CLIP (w/ respective visual backbones) on the corresponding source datasets.
+6.3 Additional Results
+In the following subsections, we provide additional results and discussions.
+6.3.1 Zero-shot performance analysis of ViT-B/32 and ViT-L/14
+In the main paper, we analyze and evaluate the zero-shot performance of ResNet-101 (RN101) [17] and ViT-B/16 [11] and conclude that such CLIP backbones are extremely sensitive, in terms of classification accuracy, to increasing severity levels of image corruption.
+This could be a major concern in situations involving real-time deployment of CLIP.
+Here, we present a similar analysis in Figure 10, using ViT-B/32 and ViT-L/14 as backbones.
+Our analysis, from the main paper, carries forward.
+To summarise, irrespective of the CLIP visual backbone, the robustness towards image corruption is limited.
+The classification performance degrades with an increase in the severity of corruption in an image.
+Method Venue
 Gaussian
@@
 CIFAR-10C
-ViT-B/16
-37.92 41.70 54.42 71.75 40.89 67.93 73.62 73.89 77.35 70.22 84.45 62.36 53.81 47.65 59.43 61.16
+ZS
+ICLR’21 35.47 39.94 43.23 69.95 41.43 64.50 70.13 70.85 72.33 66.66 81.37 64.57 59.69 48.28 56.62 59.00
 TENT
-15.49 18.28 38.12 81.59 21.73 76.32 82.35 84.62 82.19 80.60 91.83 80.55 63.52 58.57 54.71 62.03
+ICLR’21 20.09 23.45 34.47 69.85 23.01 39.79 60.35 76.83 77.49 76.07 88.88 81.38 65.35 57.01 51.19 56.35
 RoTTA
-39.17 42.90 55.41 72.18 41.32 68.02 74.01 74.38 78.01 70.80 84.80 63.19 54.62 49.32 60.15 61.89
+CVPR’23 36.55 40.91 43.99 70.03 42.45 64.52 70.08 71.23 72.68 67.31 81.92 64.99 60.33 49.40 57.11 59.57
 RPL
-15.47 17.43 40.73 81.76 20.08 69.89 82.93 84.43 83.19 81.84 91.80 79.42 64.89 54.07 54.90 61.52
+arXiv 15.89 19.08 34.04 77.84 18.72 41.22 62.39 78.17 78.86 76.31 88.83 81.15 68.98 54.19 51.91 56.51
 SAR
-47.98 53.60 60.56 74.30 47.56 73.15 76.43 77.91 79.88 75.66 86.79 71.62 58.34 62.03 64.71 67.37
-Ours
-61.13 64.09 65.76 80.51 54.96 80.65 81.94 83.04 84.19 80.84 88.95 82.15 69.16 62.68 67.64 73.85
-ViT-B/32
-35.47 39.94 43.23 69.95 41.43 64.50 70.13 70.85 72.33 66.66 81.37 64.57 59.69 48.28 56.62 59.00
-TENT
-20.09 23.45 34.47 69.85 23.01 39.79 60.35 76.83 77.49 76.07 88.88 81.38 65.35 57.01 51.19 56.35
-RoTTA
-36.55 40.91 43.99 70.03 42.45 64.52 70.08 71.23 72.68 67.31 81.92 64.99 60.33 49.40 57.11 59.57
-RPL
-15.89 19.08 34.04 77.84 18.72 41.22 62.39 78.17 78.86 76.31 88.83 81.15 68.98 54.19 51.91 56.51
-SAR
-50.28 54.12 49.65 73.08 51.98 71.17 74.65 73.73 75.22 70.99 84.25 72.08 63.93 51.57 60.32 65.13
+ICML’22 50.28 54.12 49.65 73.08 51.98 71.17 74.65 73.73 75.22 70.99 84.25 72.08 63.93 51.57 60.32 65.13
+[…]
+TPT
+NeurIPS’22 43.11 46.53 48.29 71.31 47.80 66.89 71.96 74.00 76.00 68.81 84.12 66.35 63.86 51.86 58.01 62.59
+VTE
+ECCV-W’24 47.59 50.18 53.15 71.39 53.86 67.92 72.90 76.37 76.30 70.78 83.27 61.07 69.00 58.57 61.14 64.90
+WATT-P*
+NeurIPS’24 43.64 47.1 45.97 74.98 48.04 70.42 74.74 76.1 76.86 71.85 85.15 70.36 64.6 56.0 60.37 64.41
+WATT-S*
+NeurIPS’24 56.38 58.08 52.48 78.07 58.29 76.42 79.16 78.9 79.71 76.76 87.19 76.49 70.35 64.11 65.05 70.49
 Ours
 52.39 55.99 52.54 76.79 54.04 74.90 75.79 77.67 79.10 75.31 86.33 77.34 67.41 57.06 61.29 68.26
-Table 7: Mean accuracy (%) on CIFAR-10C - TTA mean accuracy of the 15 corruptions (tasks) at a severity level of 5, using ViT-B/16 and ViT-B/32.
-We contrast our results against zero-shot backbones, TENT [46], RoTTA [47], RPL [34], and SAR [28].
-Method
-Gaussian
-Shot
-Impulse
-Defocus
-Glass
-Motion
-Zoom
-Snow
-Frost
-Fog
-Brightness
-Contrast
-Elastic
-Pixelate
-JPEG
-Mean
 CIFAR-100C
-ViT-B/16
-19.64 21.40 25.26 42.54 20.03 43.17 47.95 48.35 49.74 41.57 57.02 34.58 29.15 23.96 32.43 35.79
+ZS
+ICLR’21 16.23 17.83 17.57 39.07 17.63 38.55 43.81 42.32 43.46 39.71 50.32 29.34 28.74 22.85 29.42 31.79
 TENT
-7.60 8.21 8.33 51.81 7.95 52.45 55.34 54.16 36.17 50.92 65.63 54.51 36.52 43.99 35.81 37.96
+ICLR’21 5.53 7.64 6.85 49.60 4.47 48.45 52.35 49.77 26.77 37.50 63.05 50.53 13.89 27.00 30.80 31.61
 RoTTA
-20.65 22.22 26.17 42.48 20.26 42.90 47.88 48.75 49.92 41.86 57.00 34.52 29.27 25.08 32.88 36.12
+CVPR’23 16.63 18.25 17.78 38.62 17.76 38.38 43.52 42.39 43.41 39.37 50.60 28.85 28.89 23.50 29.65 31.84
 RPL
-6.44 7.09 7.09 52.16 11.81 52.33 55.50 54.20 38.83 51.99 66.07 54.45 36.86 42.83 39.45 38.47
+arXiv 4.50 5.80 9.61 50.26 4.43 48.88 52.61 50.27 22.36 25.34 63.36 50.31 9.10 18.65 34.53 30.00
 SAR
-25.30 27.19 32.78 47.12 23.42 47.16 51.70 51.94 52.48 48.77 61.54 44.50 32.26 33.67 38.06 41.19
-Ours
-24.91 27.73 33.66 50.11 26.27 48.49 54.85 52.35 51.62 48.38 63.27 45.21 34.74 32.38 37.31 42.09
-ViT-B/32
-16.23 17.83 17.57 39.07 17.63 38.55 43.81 42.32 43.46 39.71 50.32 29.34 28.74 22.85 29.42 31.79
-TENT
-5.53 7.64 6.85 49.60 4.47 48.45 52.35 49.77 26.77 37.50 63.05 50.53 13.89 27.00 30.80 31.61
-RoTTA
-16.63 18.25 17.78 38.62 17.76 38.38 43.52 42.39 43.41 39.37 50.60 28.85 28.89 23.50 29.65 31.84
-RPL
-4.50 5.80 9.61 50.26 4.43 48.88 52.61 50.27 22.36 25.34 63.36 50.31 9.10 18.65 34.53 30.00
-SAR
-24.63 27.14 21.25 44.57 22.98 43.95 48.40 48.01 47.76 44.85 57.76 42.11 32.69 28.02 33.08 37.81
+ICML’22 24.63 27.14 21.25 44.57 22.98 43.95 48.40 48.01 47.76 44.85 57.76 42.11 32.69 28.02 33.08 37.81
+[…]
+TPT
+NeurIPS’22 16.08 17.65 17.54 39.21 19.47 38.91 44.01 43.45 44.46 40.15 50.93 27.77 30.91 23.36 29.55 32.23
+VTE
+ECCV-W’24 16.84 18.33 18.94 39.63 22.88 39.13 43.80 44.56 44.88 39.21 49.37 28.37 34.13 26.87 30.12 33.14
+WATT-P*
+NeurIPS’24 15.55 17.02 16.16 40.25 16.16 38.74 43.6 42.49 43.51 39.39 51.17 31.85 28.35 23.94 29.74 31.86
+WATT-S*
+NeurIPS’24 16.22 17.72 16.85 41.54 17.04 39.66 44.55 43.33 44.26 40.26 52.13 33.13 29.34 24.65 30.39 32.73
 Ours
 21.35 24.71 22.32 46.26 23.07 44.64 50.12 47.23 46.88 44.92 58.55 38.52 34.56 27.73 33.19 37.60
-Table 8: Mean accuracy (%) on CIFAR-100C - TTA mean accuracy of the 15 corruptions (tasks) at a severity level of 5, using ViT-B/16 and ViT-B/32.
-We contrast our results against zero-shot backbones, TENT [46], RoTTA [47], RPL [34], and SAR [28].
-Method
-Gaussian
-Shot
-Impulse
-Defocus
-Glass
-Motion
-Zoom
-Snow
-Frost
-Fog
-Brightness
-Contrast
-Elastic
-Pixelate
-JPEG
-Mean
 ImageNet-C
-ViT-B/16
-11.18 12.54 12.04 23.36 15.18 24.50 22.58 32.32 29.88 35.88 54.08 17.20 12.72 30.96 33.26 24.51
+ZS
+ICLR’21 12.88 13.04 12.90 24.42 11.86 22.72 20.20 25.70 25.84 30.28 50.54 17.32 18.96 32.20 29.12 23.20
 TENT
-5.14 5.70 7.44 25.22 19.34 26.80 24.16 33.56 30.42 37.74 54.24 22.50 13.90 35.02 36.08 25.15
+ICLR’21 9.18 8.50 10.42 26.02 15.72 26.06 21.64 27.12 26.18 31.60 50.58 22.28 20.12 34.06 31.30 24.05
 RoTTA
-11.34 12.96 12.32 23.38 15.50 24.66 22.90 32.56 30.02 35.98 54.32 17.20 12.80 31.06 33.46 24.78
+CVPR’23 13.10 13.30 13.02 24.48 11.96 22.86 20.28 26.06 26.06 30.24 50.46 17.34 19.16 32.44 29.18 23.33
 RPL
-9.04 10.04 10.96 24.40 17.40 26.28 23.76 32.70 30.62 36.64 54.04 19.38 13.24 33.14 34.60 25.08
+arXiv 11.68 10.98 12.10 25.68 13.24 23.98 20.84 26.32 26.12 30.86 50.62 19.30 19.48 33.14 29.92 23.62
 SAR
-17.96 20.46 20.68 25.72 23.04 29.52 26.04 34.92 32.74 39.00 55.00 27.14 19.64 36.66 37.50 29.73
+ICML’22 19.82 20.36 20.92 25.78 20.40 28.34 23.10 28.12 28.38 34.74 51.10 24.60 24.38 36.54 34.40 28.07
+[…]
+TPT
+NeurIPS’22 12.04 12.64 12.52 25.38 12.28 22.68 20.78 26.36 26.64 30.78 51.02 16.50 19.90 33.62 30.62 23.58
+VTE
+ECCV-W’24 11.96 12.32 13.44 25.06 11.70 22.58 22.40 27.38 27.02 32.28 51.52 16.84 19.94 34.80 32.82 24.14
+StatA ($\gamma$=0.1)
+CVPR’25 11.57 12.28 11.74 21.91 10.91 20.43 18.87 24.04 25.0 29.12 49.21 16.02 18.87 29.15 26.72 21.73
+StatA ($\gamma$=-1)
+CVPR’25 12.52 13.10 12.74 22.43 11.39 21.17 19.61 24.37 25.60 29.44 49.89 17.71 19.21 29.87 27.52 22.44
 Ours
-19.32 21.38 19.60 26.58 21.94 30.88 29.02 36.48 32.00 40.98 56.72 26.14 23.74 37.68 38.34 30.72
-ViT-B/32
-12.88 13.04 12.90 24.42 11.86 22.72 20.20 25.70 25.84 30.28 50.54 17.32 18.96 32.20 29.12 23.20
-TENT
-9.18 8.50 10.42 26.02 15.72 26.06 21.64 27.12 26.18 31.60 50.58 22.28 20.12 34.06 31.30 24.05
-RoTTA
-13.10 13.30 13.02 24.48 11.96 22.86 20.28 26.06 26.06 30.24 50.46 17.34 19.16 32.44 29.18 23.33
-RPL
-11.68 10.98 12.10 25.68 13.24 23.98 20.84 26.32 26.12 30.86 50.62 19.30 19.48 33.14 29.92 23.62
-SAR
-19.82 20.36 20.92 25.78 20.40 28.34 23.10 28.12 28.38 34.74 51.10 24.60 24.38 36.54 34.40 28.07
-Ours
-16.84 18.20 16.10 25.04 20.90 28.90 25.24 29.42 27.18 36.02 50.18 17.66 27.58 36.20 35.42 27.39
-Table 9: Mean accuracy (%) on ImageNet-C - TTA mean accuracy of the 15 corruptions (tasks) at a severity level of 5, using ViT-B/16 and ViT-B/32.
-We contrast our results against zero-shot backbones, TENT [46], RoTTA [47], RPL [34], and SAR [28].
-6.2 Implementation Details
-In this section, we summarize the implementation details of all the baseline methods that have been mentioned in the main paper, including ours.
-We build our approach on the open-source code base  [^3] https://github.com/mariodoebler/test-time-adaptation/tree/main.
-a standard TTA benchmark codebase, that also houses the hyperparameters and training details of all the prior TTA methods.
-CLIP-like models are used as provided by OpenCLIP.
-Only the vision encoder is updated for online TTA methods [46, 47, 34, 28] adopted for CLIP.
-BAT-CLIP (Ours): For domain-specific test adaptation, we conducted experiments using ViT-B/16 and ViT-B/32 [10] as the vision backbones.
-For CIFAR-10C, both the vision encoder ($f_{vis}$) and text encoder ($f_{txt}$) were updated using the AdamW optimizer with a learning rate of 10-3.
-Similarly, for CIFAR-100C and ImageNet-C, we employed the Adam optimizer and AdamW optimizer, respectively, with a learning rate of 5$\times$10-4.
-The batch size $\mathcal{B}$ used was set to 200 for CIFAR-10C and CIFAR-100C, and 64 for ImageNet-C.
-Throughout, the prompt template is fixed to “a photo of a ¡CLS¿.”.
-TPT [38]: For each test image, 63 augmentations are generated based on random resized crops, yielding a batch of 64 images, in addition to the original test image.
-The prompt/context vectors are initialized based on “a photo of a ¡CLS¿.” and tokenized using pre-trained CLIP weights.
-The confidence threshold is set to 10% i.e., the marginal entropy over the 10% confident samples is minimized.
-For all the datasets, we follow their core implementation and optimize the prompt vectors using an AdamW optimizer with a learning rate of 5$\times$10-3.
-VTE [9]: In VTE, an ensemble of different prompt templates is considered based on the idea of [31].
-An example of templates includes “a photo of a ¡CLS¿.”, “a sketch of a ¡CLS¿.”, “a painting of a ¡CLS¿.”, etc.
-The prompt templates are then averaged.
-On the vision side, similar to TPT [38], a batch of random augmentation is created for a test image with no model updates.
-TENT [46]: We follow all the hyperparameters that TENT provides in their official implementation  [^4] https://github.com/DequanWang/tent.
-To update the vision encoder, we use Adam as the optimizer with a learning rate of 10-3 for CIFAR-10C and CIFAR-100C.
-For ImageNet-C, we update using SGD with a learning rate of 25$\times$10-5.
-RoTTA [47]: For fairness, the batch sizes are set to 200 for the CIFAR datasets and 64 for ImageNet-C.
-The vision encoder is updated based on the Adam rule with a learning rate of 10-3.
-The capacity of the memory bank is set to 64, for all the datasets.
-Following the notations in the paper, $\alpha$ = 0.05, $\delta$ = 0.1, $\nu$ = 0.001, $\lambda_{t}$ and $\lambda_{u}$ = 1.0.
-We implement the details exactly as described in their main paper.
-RPL [34]: We use an Adam optimizer with a learning rate of 10-3 for CIFAR-10C and CIFAR-100C.
-For ImageNet-C, the update rule is SGD with a learning rate of 5$\times$10-4.
-To compute the generalized cross-entropy loss, q is set to 0.8 for all the datasets.
-SAR [28]: The training details/hyperparameters for SAR are the same as RPL [34] for CIFAR-10 and CIFAR-100.
-For ImageNet-C, the learning rate is set to 25$\times$10-5 with an SGD update rule.
-The entropy threshold $E_{0}$ is 0.4xln(C), where C is the number of classes.
-$\rho$ is set to a default of 0.05.
-The moving average factor is 0.9 for $e_{m}$ and $e_{0}$ is set to 0.2.
-All parameters are the same as in [28].
-Figure 9: Task-wise mean accuracy (%) of zero-shot CLIP across different corruption severity levels.
-[Top]: ViT-B/32 backbone.
-[Bottom]: ViT-L/14 backbone.
-The dashed lines indicate the performance of zero-shot CLIP (w/ respective visual backbones) on the corresponding source datasets.
-6.3 Additional Results
-In the following subsections, we provide additional results and discussions.
-6.3.1 Zero-shot performance analysis of ViT-B/32 and ViT-L/14
-In the main paper, we analyze and evaluate the zero-shot performance of ResNet-101 (RN101) [14] and ViT-B-16 [10] and conclude that such CLIP backbones are extremely sensitive, in terms of classification accuracy, to increasing severity levels of image corruption.
-This could be a major concern in situations involving real-time deployment of CLIP.
-Here, we present a similar analysis in Figure 9, using ViT-B/32 and ViT-L/14 as backbones.
-Our analysis, from the main paper, carries forward.
-To summarise, irrespective of the CLIP visual backbone, the robustness towards image corruption is limited.
-The classification performance degrades with an increase in the severity of corruption in an image.
-Method
-Gaussian
-Shot
-Impulse
-Defocus
-Glass
-Motion
-Zoom
-Snow
-Frost
-Fog
-Brightness
-Contrast
-Elastic
-Pixelate
-JPEG
-Zero-Shot
-ViT-B/16
-90.1
-Ours
-84.51 84.29 88.69 88.48 86.44 86.46 87.04 91.38 91.01 90.22 90.87 88.12 88.13 74.74 87.48 87.19 (mean)
-ViT-B/32
-88.3
-Ours
-67.41 68.28 84.23 80.50 75.37 79.75 78.55 87.67 86.36 85.83 90.04 80.89 81.56 82.74 82.36 80.77 (mean)
-Table 10: Zero-shot performance on CIFAR10 (source) after adaptation of BAT-CLIP on a task.
-Method
-Gaussian
-Shot
-Impulse
-Defocus
-Glass
-Motion
-Zoom
-Snow
-Frost
-Fog
-Brightness
-Contrast
-Elastic
-Pixelate
-JPEG
-Zero-Shot
-ViT-B/16
-66.6
-Ours
-67.09 67.12 67.08 70.31 63.70 66.83 70.12 70.10 68.19 69.55 71.05 66.49 65.13 59.98 67.60 67.36 (mean)
-ViT-B/32
-62.3
-Ours
-45.99 46.90 60.86 63.62 57.73 59.59 61.66 66.21 61.50 63.61 66.39 57.05 62.13 62.82 65.37 60.09 (mean)
-Table 11: Zero-shot performance on CIFAR100 (source) after adaptation of BAT-CLIP on a task.
-6.3.2 Results against prior TTA methods
-We adopt prior TTA methods [46, 47, 34, 28] for CLIP and present the fine-grained results i.e., the task-wise mean accuracy to demonstrate the efficacy of our method (the higher the better).
-The results are reported in Tables 7, 8, and 9.
-Across all the backbones and datasets, we see that our method BAT-CLIP, achieves the best or comparable performance against all the baseline TTA approaches adopted for CLIP.
+16.84 18.20 16.10 25.04 20.90 28.90 25.24 29.42 27.18 36.02 50.18 17.66 27.68 36.20 35.42 27.39
+Table 5: Mean accuracy (%) on CIFAR-10C, CIFAR-100C, and ImageNet-C - TTA mean accuracy of the 15 corruptions (tasks) at a severity level of 5, using ViT-B/32.
+6.3.2 Online TTA results using a ViT-B/32 backbone
+In the main paper, we had presented the online TTA results on CIFAR-10C, CIFAR-100C, and ImageNet-C, using a ViT-B/16 backbone.
+In Table 5, we provide results using a ViT-B/32 backbone.
+Across all the datasets, we see that BATCLIP achieves the best or comparable performance against all the baseline approaches.
 6.3.3 Detailed results from the loss ablation study
-In the main paper, we provide ablation of loss components and their combinations i.e., the mean accuracy across all the tasks for ViT-B/16 and ViT-B/32 on the benchmarks corruption datasets.
+In the main paper, we provide ablation of loss components and their combinations i.e., the mean accuracy across all the tasks for ViT-B/16 on the benchmark corruption datasets.
 Here, we provide additional task-wise accuracy in Table 6.
@@
 Table 6: Task-wise loss ablation results (accuracy) on CIFAR-10C, CIFAR-100C, and ImageNet-C.
+Method
+Gaussian
+Shot
+Impulse
+Defocus
+Glass
+Motion
+Zoom
+Snow
+Frost
+Fog
+Brightness
+Contrast
+Elastic
+Pixelate
+JPEG
+Zero-Shot
+ViT-B/16
+90.1
+Ours
+84.51 84.29 88.69 88.48 86.44 86.46 87.04 91.38 91.01 90.22 90.87 88.12 88.13 74.74 87.48 87.19 (mean)
+ViT-B/32
+88.3
+Ours
+67.41 68.28 84.23 80.50 75.37 79.75 78.55 87.67 86.36 85.83 90.04 80.89 81.56 82.74 82.36 80.77 (mean)
+Table 7: Zero-shot performance on CIFAR10 (source) after adaptation of BATCLIP on a task.
+Method
+Gaussian
+Shot
+Impulse
+Defocus
+Glass
+Motion
+Zoom
+Snow
+Frost
+Fog
+Brightness
+Contrast
+Elastic
+Pixelate
+JPEG
+Zero-Shot
+ViT-B/16
+66.6
+Ours
+67.09 67.12 67.08 70.31 63.70 66.83 70.12 70.10 68.19 69.55 71.05 66.49 65.13 59.98 67.60 67.36 (mean)
+ViT-B/32
+62.3
+Ours
+45.99 46.90 60.86 63.62 57.73 59.59 61.66 66.21 61.50 63.61 66.39 57.05 62.13 62.82 65.37 60.09 (mean)
+Table 8: Zero-shot performance on CIFAR100 (source) after adaptation of BATCLIP on a task.
+6.3.4 Effect of different prompt templates
+In Table , we show results with “relevant” prompt templates to show the independence of such prompt selection, at test-time.
+As seen, the performance gain over zero-shot ViT-B/32 is fairly large for all the prompt templates.
+Though TPT [45] fine-tunes a pre-trained prompt on each image, and VTE [10] uses an ensemble of prompts, our method is agnostic to the prompt template being used, making it favorable for real-time usage.
+In all of our prior experiments, we use a generic prompt template “a photo of a $<$CLS$>$.” for all of the datasets and methods.
+Here, we replace this with “relevant” prompt templates to show the independence of such a prompt selection, at test-time, and report the results in Table 9.
+As seen, the performance gain over zero-shot ViT-B/32 is fairly large for all the prompt templates.
+Though TPT [45] fine-tunes a pre-trained prompt on each test image, and VTE [10] uses an ensemble of prompts, our method is agnostic to the prompt template being used, making it favorable for real-time deployment.
+Prompt Template CIFAR-10C CIFAR-100C ImageNet-C
+a low contrast photo of a $<$CLS$>$.
+68.53 (+7.81) 37.09 (+4.97) 27.31 (+3.71)
+a blurry photo of a $<$CLS$>$.
+68.84 (+10.96) 36.80 (+5.33) 26.92 (+3.52)
+a photo of a big $<$CLS$>$.
+67.49 (+10.10) 35.79 (+4.87) 25.64 (+3.29)
+Table 9: Prompt template selection. + denotes the accuracy gain over zero-shot ViT-B/32.
+6.3.5 Impact of bimodal adaptation
+To show the effectiveness of bimodal adaptation, we ablate $\mathcal{L}_{pm}$, which is responsible for updating the text encoder $f_{txt}$.
+We report the results in Table 10.
+We see a drop in accuracy when $f_{txt}$ is “frozen” i.e., when $\mathcal{L}_{pm}$ isn’t used, necessitating the need for bimodal adaptation of CLIP encoders.
+Dataset $f_{vis}$ update $f_{txt}$ update Ours
+CIFAR-10C 72.58
+73.85
+CIFAR-100C 41.00
+42.09
+ImageNet-C 29.88
+30.72
+Table 10: Ablation on $\mathcal{L}_{pm}$ to demonstrate the need of bimodal adaptation of CLIP encoders - using a ViT-B/16 backbone.
 6.3.6 Post-adaptation results on source test sets
@@
 Given that CLIP has been adapted to a specific corruption task, will the zero-shot generalization still hold back on its source test set?
-In this crucial experiment, we challenge our BAT-CLIP, and evaluate its zero-shot generalization performance back on the source test set, to check the preservation of pre-trained CLIP Knowledge.
+In this crucial experiment, we challenge our BATCLIP, and evaluate its zero-shot generalization performance back on the source test set, to check the preservation of pre-trained CLIP Knowledge.
 After the adaptation of CLIP on each corruption task, we report the adapted model’s zero-shot performance on its corresponding source test set.
-We report results for CIFAR-10C and CIFAR-100C in Tables 10 and 11, using ViT-B/16 and ViT-B/32 backbones.
-For all of the results, we use the prompt template “a photo of a ¡CLS¿.”.
+We report results for CIFAR-10C and CIFAR-100C in Tables 7 and 8, using ViT-B/16 and ViT-B/32 backbones.
+For all of the results, we use the prompt template “a photo of a $<$CLS$>$.”.
 As an example, for CIFAR-10C, upon adaptation of CLIP to Gaussian noise following our approach, we report the adapted model’s zero-shot accuracy on its source test set - CIFAR10 test set.
@@
 Overall, ViT-B/16 visual backbones preserve larger amounts of CLIP pre-trained knowledge.
-This proves the effectiveness of our method BAT-CLIP, on average.
+This proves the effectiveness of our method BATCLIP, on average.
 6.3.7 t-SNE visualizations of CIFAR-10C and CIFAR-100C
 In this section, we provide illustrations of task-wise t-SNE [49] plots for CIFAR-10C and CIFAR-100C and compare them against zero-shot ViT-B/16.
-The results are in Figures 10, 11, 12, and 13.
-Across all corruptions/tasks, BAT-CLIP learns strong discriminative visual features with a strong image-text alignment and class-level separation.
+The results are in Figures 11, 12, 13, and 14.
+Across all corruptions/tasks, BATCLIP learns strong discriminative visual features with a strong image-text alignment and class-level separation.
 ImageNet-C has 1000 classes, so, we do not provide t-SNE plots to avoid complications.
@@
 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Figure 10: BAT-CLIP (w/ ViT-B/16): The t-SNE plots show visual ($\circ$) and text ($\bigstar$) features for CIFAR-10C.
+Figure 11: BATCLIP (w/ ViT-B/16): The t-SNE plots show visual ($\circ$) and text ($\bigstar$) features for CIFAR-10C.
 Gaussian Shot Impulse Defocus Glass
@@
 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Figure 12: BAT-CLIP (w/ ViT-B/16): The t-SNE plots show visual ($\circ$) and text ($\bigstar$) features for CIFAR-100C.
+Figure 13: BATCLIP (w/ ViT-B/16): The t-SNE plots show visual ($\circ$) and text ($\bigstar$) features for CIFAR-100C.
 Gaussian Shot Impulse Defocus Glass

### v3

@@
 Furthermore, we evaluate our proposed TTA approach on various domain generalization datasets to demonstrate its generalization capabilities.
+Our code is available at https://github.com/sarthaxxxxx/BATCLIP.
 1 Introduction
@@
 To generalize our method beyond common corruptions, we evaluate on various domain generalization datasets [15]-OfficeHome [52], PACS [26], VLCS [13], and Terra Incognita [4].
-Overall, our approach achieves state-of-the-art performance for CLIP adaptation at test-time on various domain shifts.
+Overall, our approach achieves SOTA performance for CLIP adaptation at test-time on various domain shifts (see Supplementary).
 2 Related Works
@@
 $\text{sim}(v,z)$ $=\frac{v^{T}\cdot z}{||v||_{2}\cdot||z||_{2}}$
-where $\text{sim}(\cdot)$ is the cosine similarity and $\tau$ is the softmax temperature from CLIP’s pre-training stage.
-The text features $\mathcal{Z}$, in zero-shot evaluation, are always pre-computed.
-We also draw a comparison with CLIP performance on respective source test sets and follow this implementation  [^1] https://github.com/LAION-AI/CLIP_benchmark.
-Throughout this paper, the same TTA problem setup is used, unless mentioned otherwise.
-This TTA setup differs from the TTA setups of [36, 15, 34].
+[…]
 Figure 2: Task-wise mean accuracy (%) of zero-shot CLIP across corruption severity levels.
@@
 Similarly, the performance of the text encoder $f_{txt}$ is invariant to the different text prompt templates and is also impractical to tune at test-time.
-Therefore, for effective adaptation, both the image and text encoders of CLIP need to be adapted to the incoming domain of test batches.
-We illustrate our proposed framework, BATCLIP, in Figure 4.
+Therefore, for effective adaptation, both the image and text encoders of CLIP need to be adapted to the incoming domain of test batches to focus on achieving strong image-text alignment and feature separability.
+We illustrate BATCLIP in Figure 4.
 From unimodal adaptation to bimodal adaptation.
@@
 Motivated by these works, we update the LayerNorm parameters of both CLIP encoders.
-This constitutes updating $\sim 0.044$% of all parameters.
+This constitutes updating 0.044$\sim$ 0.044∼ 0.044% of all parameters.
 For every new task/corruption, we reset the model parameters following TENT [54] i.e., we perform a single domain TTA.
@@
 We compare our approach with zero-shot CLIP (ZS), state-of-the-art online TTA methods adopted for CLIP following [11] i.e., TENT [54], RoTTA [55], RPL [42], and SAR [36].
-We update only the vision encoder $f_{vis}$ for such a setup.
+We update only the vision encoder $f_{vis}$ as in the baselines.
 We benchmark against other online TTA methods using CLIP - TPT [46] and VTE [11].
@@
 We mention the details of each TTA method in the Supplementary.
-Implementation Details.
-We query the text encoder $f_{txt}$ with a general prompt template “a photo of a $<$CLS$>$.” for all the datasets, as motivated earlier.
-We optimize both the vision encoder $f_{vis}$ and text encoder $f_{txt}$.
-For CIFAR-10C, we use an AdamW optimizer at a learning rate of $10^{-3}$.
-Similarly, for CIFAR-100C and ImageNet-C, Adam and AdamW optimizers are respectively used, at a fixed learning rate of 5$\times$$10^{-4}$, with the model being reset after each task.
-For fairness to all the baselines, the batch sizes are set to 200, 200, and 64 for the datasets, following various TTA benchmarks, at a corruption severity level of 5 for each task.
-For WATT-P* and WATT-S*, we use the same templates and learning rates as mentioned in [36].
-For StatA, the class correlation in a batch is controlled by a Dirichlet distribution ($\gamma$).
-We report results for $\gamma$ = 0.1 (low correlation) and -1 (sequential).
-All experiments are run on a single NVIDIA RTX A5000 GPU using ViT-B/16 as the visual backbone.
-Results on ViT-B/32 are reported in the Supplementary.
+[…]
 Method Venue
@@
 ICML’22 47.98 53.60 60.56 74.30 47.56 73.15 76.43 77.91 79.88 75.66 86.79 71.62 58.34 62.03 64.71 67.37
-[…]
 TPT
@@
 ICML’22 25.30 27.19 32.78 47.12 23.42 47.16 51.70 51.94 52.48 48.77 61.54 44.50 32.26 33.67 38.06 41.19
-[…]
 TPT
@@
 ICML’22 17.96 20.46 20.68 25.72 23.04 29.52 26.04 34.92 32.74 39.00 55.00 27.14 19.64 36.66 37.50 29.73
-[…]
 TPT
@@
 The largest performance improvements come on the Terra Incognita dataset.
+Large style variations in OfficeHome and PACS make it difficult for CLIP to generate effective text features, with a single step, from our default prompt.
 The implementation details are described in the Supplementary.
@@
 Figure 6: Analysis on increasing # iterations: Mean accuracy for # iterations, on each test batch, using a ViT-B/16 backbone.
+Backbone (ViT-B/16) CIFAR-10C CIFAR-100C ImageNet-C
+$\mathcal{L}_{ent}$ 60.65 38.17 24.03
+$\mathcal{L}_{pm}$ 61.23 36.63 22.83
+$\mathcal{L}_{sp}$ 73.16 41.36 30.05
+$\mathcal{L}_{ent}$ + $\mathcal{L}_{pm}$ 62.60 39.32 25.21
+$\mathcal{L}_{ent}$ + $\mathcal{L}_{sp}$ 72.69 41.84 30.08
+[…]
+[…]
+$\mathcal{L}_{ent}$ + $0.3(\mathcal{L}_{pm}$+$\mathcal{L}_{sp}$) 73.31 41.93 30.40
+$\mathcal{L}_{ent}$ + $0.7(\mathcal{L}_{pm}$ +$\mathcal{L}_{sp}$) 73.85 42.04 30.61
+$\mathcal{L}_{ent}$ + $\mathcal{L}_{pm}$ + $\mathcal{L}_{sp}$ 73.85 42.09 30.72
+Table 3: Ablation on loss components - Mean accuracy (in %).
 Adaptation for multiple iterations.
@@
 In the Supplementary, we report the post-adaptation results back on the source test sets - CIFAR10 and CIFAR100 to evaluate catastrophic forgetting.
-Our approach outperforms zero-shot CLIP, with a smaller decline in mean accuracy—especially on CIFAR-10C and ImageNet-C—showing that our loss components improve robustness under challenging conditions.
-Backbone (ViT-B/16) CIFAR-10C CIFAR-100C ImageNet-C
-$\mathcal{L}_{ent}$ 60.65 38.17 24.03
-$\mathcal{L}_{pm}$ 61.23 36.63 22.83
-$\mathcal{L}_{sp}$ 73.16 41.36 30.05
-$\mathcal{L}_{ent}$ + $\mathcal{L}_{pm}$ 62.60 39.32 25.21
-$\mathcal{L}_{ent}$ + 0.1 * ($\mathcal{L}_{pm}$ + $\mathcal{L}_{sp}$) 71.90 41.88 30.24
-$\mathcal{L}_{ent}$ + 0.5 * ($\mathcal{L}_{pm}$ + $\mathcal{L}_{sp}$) 73.70 41.98 30.44
-$\mathcal{L}_{ent}$ + $\mathcal{L}_{pm}$ + $\mathcal{L}_{sp}$ 73.85 42.09 30.72
-Table 3: Ablation on loss components - Mean accuracy (in %).
+Our approach outperforms zero-shot CLIP, with a smaller decline in mean accuracy, especially on CIFAR-10C and ImageNet-C, showing that our losses improve robustness.
+under challenging conditions.
 Ablation on loss components.
-In Table 3, we ablate the loss components by taking their different combinations using ViT-B/16.
-The task-wise results are reported in the Supplementary.
-We observe that the addition of $\mathcal{L}_{pm}$ and $\mathcal{L}_{sp}$ largely improve model performance than entropy minimization via $\mathcal{L}_{ent}$.
+In Table 3, we ablate the loss components using ViT-B/16 (see more in the Supplementary).
+$\mathcal{L}_{pm}$ (row 4) and $\mathcal{L}_{sp}$ (row 5) improve model performance than entropy minimization via $\mathcal{L}_{ent}$.
 In fact, the addition of $\mathcal{L}_{sp}$ brings larger improvements, proving that $f_{vis}$ produces discriminative features.
-Interestingly, using just $\mathcal{L}_{ent}$ achieves better or comparable accuracy than zero-shot CLIP, as the encoders work in synergy while adapting to the domain.
+Interestingly, using only $\mathcal{L}_{ent}$ achieves comparable or better accuracy than zero-shot CLIP, with fully hyperparameter-free objectives.
 Refer to caption Refer to caption
@@
 (a) Mean acc. w/. a ViT-L/14.
-Dataset TENT RPL RoTTA SAR Ours
-ImageNet-C 40.62 40.00 39.76 42.10 42.84
-(b) TTA baselines vs ours w/ ViT-L/14.
-Table 4: Ablation on a larger backbone (ViT-L/14) - Mean accuracy (in %)
+TTA Method ImageNet-C
+TENT 40.62
+RPL 40.00
+RoTTA 39.76
+SAR 42.10
+Ours 42.84
+(b) Results on ImageNet-C.
+Table 4: Ablation on ViT-L/14 - Mean accuracy (in %).
 Extension to larger backbones.
-We run BATCLIP for a larger backbone like ViT-L/14 and report the results in Table 4.
+We use ViT-L/14 and report the results in Table 4.
 The key advantage lies in the projection matching loss ($\mathcal{L}_{pm}$) which updates the text encoder to produce more image-aware text features—an aspect missing in other state-of-the-art TTA baselines.
 Computational efficiency.
-Due to an additional text encoder update, BATCLIP updates $\sim 0.044$% of all CLIP parameters, with a ViT-B/16 backbone, as opposed to $\sim 0.026$% in other TTA baselines.
+Due to an additional text encoder update, BATCLIP updates 0.044$\sim$ 0.044∼ 0.044% of all CLIP parameters, with a ViT-B/16 backbone, as opposed to 0.026$\sim$ 0.026∼ 0.026% in other TTA baselines.
 Though the increase is marginal, this leads to superior performances as adaptation of the text encoder produces input-aware CLIP text features, aligning better with the image features.
@@
 Empirical studies, including ablations, show robust performance gains over all baselines on benchmark corruption and domain generalization datasets.
+7 Acknowledgement
+This project was supported by a grant from UT Dallas and an NVIDIA Academic Grant Program Award.
 Supplementary Material
@@
 Subsection 7.3.5 discusses the ablation of $\mathcal{L}_{pm}$ which is responsible for updating the text encoder.
-    • In subsections 6.3.3 and 6.3.6, we present a detailed loss ablation study and the post-adaptation zero-shot generalization on source test sets, respectively — essentially evaluating catastrophic forgetting of BATCLIP.
+    • In subsections 7.3.3 and 7.3.8, we present a detailed loss ablation study and the post-adaptation zero-shot generalization on source test sets, respectively — essentially evaluating catastrophic forgetting of BATCLIP.
     • Lastly, we include task-wise t-SNE visualizations in subsection 7.3.9 for CIFAR-10C and CIFAR-100C, comparing our method against zero-shot CLIP (ViT-B/16), to illustrate the effectiveness of BATCLIP.
+Gaussian Shot Impulse Defocus Glass
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Motion Zoom Snow Frost Fog
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Brightness Contrast Elastic Pixelate JPEG
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Figure 8: We provide visualizations of an image from ImageNet-C [19] for different corruption types, at an image severity level of 5.
 7.1 Datasets
@@
 We urge the readers to check out [19] for further inspection.
-Gaussian Shot Impulse Defocus Glass
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Motion Zoom Snow Frost Fog
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Brightness Contrast Elastic Pixelate JPEG
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Figure 8: We provide visualizations of an image from ImageNet-C [18] for different corruption types, at an image severity level of 5.
 Figure 9: Comparison of classification predictions across various methods (Zero-shot CLIP (ViT-B/16), VTE [11], TPT [46], and Ours) on ImageNet-C samples with Gaussian noise.
@@
 ICML’22 50.28 54.12 49.65 73.08 51.98 71.17 74.65 73.73 75.22 70.99 84.25 72.08 63.93 51.57 60.32 65.13
-[…]
 TPT
@@
 ICML’22 24.63 27.14 21.25 44.57 22.98 43.95 48.40 48.01 47.76 44.85 57.76 42.11 32.69 28.02 33.08 37.81
-[…]
 TPT
@@
 ICML’22 19.82 20.36 20.92 25.78 20.40 28.34 23.10 28.12 28.38 34.74 51.10 24.60 24.38 36.54 34.40 28.07
-[…]
 TPT
@@
 Table 10: Ablation on $\mathcal{L}_{pm}$ to demonstrate the need of bimodal adaptation of CLIP encoders - using a ViT-B/16 backbone.
+Camera Sensor Condition ZS TENT SAR Ours
+Auto-Exposure Light on 50.90 51.62 52.60 53.82
+Light off 46.96 47.44 47.88 49.32
+Manual Light on 60.44 60.64 60.30 61.26
+Light off 60.76 60.99 59.62 61.61
+Table 11: Online TTA experiments on ImageNet-ES [2]. Mean accuracy (in %).
+SigLIP Clean (ImageNet) Source TENT SAR BATCLIP
+ImageNet-C 82.00 35.44 37.58 39.62 40.10
+Table 12: Online TTA experiments on ImageNet-C with SigLIP [58].
+7.3.6 BATCLIP for other vision-language models
+We employ SigLIP [58], a recent pre-trained vision-language model with 877 million parameters, for our online TTA experiments on ImageNet-C.
+We use a batch size of 8 and a learning rate of $5\times 10^{-5}$ with the AdamW optimizer.
+Despite the small batch size, which could impact prototypes, BATCLIP achieves better performance, as shown in Table 12.
+7.3.7 Results on distribution shifts caused by lighting conditions, camera types, or object scales
+To evaluate BATCLIP across a wide spectrum of shifts, we extend our setup to datasets exhibiting variations due to lighting conditions, camera types, and object scales.
+We conduct experiments on ImageNet-ES [2], which introduces significant variations in lighting and camera sensor settings (e.g., ISO, shutter speed).
+While more details can be found in the original work, but ImageNet-ES introduces wide variations in lighting conditions and camera sensor factors (ISO, shutter speed, etc.).
+We report the results in Table 11 using a ViT-B/16 backbone.
+To be noted that, “Auto-Exposure” has 5 tasks for each condition, while “Manual” has 27.
+On average, BATCLIP outperforms all the reported baselines.
 7.3.8 Post-adaptation results on source test sets